### v1

# On the isomorphism problem for ultraproducts of $\mathrm{C}^{*}$-algebras in continuous model theory

Abstract
In classical model theory, the Keisler-Shelah theorem establishes a fundamental connection between the elementary equivalence of structures and the isomorphism of their ultrapowers. Motivated by this, one may ask whether an analogous relationship holds in the framework of continuous model theory, which naturally encompasses metric structures such as $\mathrm{C}^{\ast}$-algebras. In this paper, we investigate the isomorphism problem for ultraproducts of operator algebras from a model-theoretic perspective. We prove that, assuming the negation of the continuum hypothesis, there exist two elementarily equivalent infinite-dimensional unital $\mathrm{C}^{\ast}$-algebras $A$ and $B$ of size $\leq\mathfrak{c}$ such that for all non-principal ultrafilters $\mathcal{U},\mathcal{V}$ on $\omega$, the ultrapowers $A^{\mathcal{U}}$ and $B^{\mathcal{V}}$ are not isomorphic. This result provides a continuous analogue of certain classical theorems concerning ultraproducts and demonstrates that the model-theoretic behavior of $\mathrm{C}^{\ast}$-algebras is closely related to set-theoretic principles such as the continuum hypothesis.
1 Introduction
The ultraproduct construction is a significant object of study in both model theory and operator algebras. The model-theoretic notion of ultraproducts was introduced by Łoś in [Los:1955] in the 1950’s, and Robinson made use of it to develop nonstandard analysis (cf. [Robinson:1966]). Around the same time, an ultraproduct-like construction appeared in the context of operator algebras in the work of Kaplansky and Wright [Wright:1954]. Subsequently, Sakai introduced the ultraproduct of $\mathrm{II}_{1}$ factors in [Sakai:1962], and McDuff [McDuff:1970] discovered its significance. Later, Connes [Connes:1976] and Kirchberg [Kirchberg:1995] applied this construction to the classification theory of operator algebras. However, no essential connection between these two perspectives seems to have been recognized at the time.
In operator algebra theory, ultraproducts over countable index sets are mainly considered, and the isomorphisms between such ultraproducts have been studied. In particular, Ge and Hadwin [GH:2001, Corollary 3.4] established the following result:
Proposition 1.1.
Assume that the continuum hypothesis holds. Let $A$ be a $\mathrm{C}^{\ast}$}-algebra with cardinality $\leq{\mathfrak{c}}$. Then, for all nonprincipal ultrafilters $\mathcal{U},\mathcal{V}$ on $\omega$, $A^{\mathcal{U}}$ and $A^{\mathcal{V}}$ are isomorphic. Moreover, if $A$ is separable, there is an isomorphism from $A^{\mathcal{U}}$ to $A^{\mathcal{V}}$ that fixes the elements of $A$. Thus the relative commutants $A^{\prime}\cap A^{\mathcal{U}}$ and $A^{\prime}\cap A^{\mathcal{V}}$ are isomorphic.
This result was given as an analogue of Keisler’s results on the saturation of ultrapowers in classical model theory. (cf. [Keisler:1961])
In recent years, continuous model theory, a generalization of classical model theory to the setting of metric structures, has gained increasing attention as a powerful tool for analyzing ultraproducts of operator algebras. The previous proposition can be proved within a model-theoretic framework using continuous model theory. Moreover, it is shown that, for all infinite-dimensional $\mathrm{C}^{\ast}$}-algebra AA, every countable ultrapowers of $A$ are isomorphic if and only if the continuum hypothesis holds (cf. [FHS1:2013, Theorem 5.1]).
In model theory, the isomorphism of ultraproducts of two structures is known to be closely related to a property called elementary equivalence. One of the earliest results in this direction was given by Keisler [Keisler:1961], which states that, under the assumption of the continuum hypothesis, for any $\mathcal{L}$-theory $\mathrm{T}$ in a countable language $\mathcal{L}$ and any two models $M$ and $N$ of $\mathrm{T}$ of size $\leq{\mathfrak{c}}$, $M$ and $N$ are elementarily equivalent if and only if they have isomorphic ultrapowers with respect to an ultrafilter on $\omega$. Furthermore, this result was generalized by Shelah [Shelah:1971], who weakened the assumption on the ultrafilter from one on $\omega$ to an ultrafilter on an arbitrary index set, and showed that the theorem holds even in ZFC. This is known as the Keisler-Shelah theorem. Moreover, it is also known to hold in the framework of continuous model theory - see [MTFMS, Theorem 5.7]. Thus, the following question naturally arises.
Question 1.2.
Assume that the continuum hypothesis fails. For which $\mathcal{L}$-theories $\mathrm{T}$ in a countable language $\mathcal{L}$, do there exist models $M$ and $N$ of $\mathrm{T}$ of size $\leq{\mathfrak{c}}$ (or $\aleph_{0}$) such that $M$ and $N$ are elementarily equivalent, but $M^{\mathcal{U}}$ and $N^{\mathcal{V}}$ are not isomorphic for all nonprincipal ultrafilters $\mathcal{U},\mathcal{V}$ on $\omega$?
In the classical case, several necessary conditions on the theory $\mathrm{T}$, as well as on concrete examples, have been established for this problem (cf. [GS1:2023], [GS2:2023], [Tsuboi:2022] and [Goto:2024]). Moreover, this problem can be naturally extended to continuous model theory, particularly in the setting of operator algebras. For instance, in [FHS3:2014, §4], this problem is discussed in the case of $\mathrm{II}_{1}$ factors. In this paper, we discuss the problem in the setting of $\mathrm{C}^{\ast}$}-algebras. The following is our main result:
Theorem 1.3.
Assume that the continuum hypothesis fails. There exists a $\mathrm{C}^{\ast}$}-algebra $M$ satisfying the following. We can find models $A$ and $B$ of the theory of $M$ with density characters $\leq{\mathfrak{c}}$ such that $A^{\mathcal{U}}$ and $B^{\mathcal{V}}$ are not isomorphic for all nonprincipal ultrafilters $\mathcal{U},\mathcal{V}$ on $\omega$.
This result contrasts with the case under the assumption of CH in section 1. For the precise conditions on $M$, see section 5.
Finally, we explain the overall outline of this paper. In Section 2, we present the formalization of continuous model theory and introduce its basic concepts. We then explain continuous analogues and proofs of several fundamental results, such as the Tarski-Vaught test and the downward Löwenheim-Skolem-Tarski theorem. In Section 3, we describe the construction of ultraproducts in continuous model theory and discuss Łoś’ theorem, together with its applications to the compactness theorem and the notion of saturation. In Section 4, we study the concept of saturation in more depth, provide a model-theoretic proof of section 1, and present its extension using the notion of types. In Section 5, we state and prove section 1, along with concrete examples of the results. Finally, in Section 6, we discuss a topological application of the results in the previous section and propose several related open problems as directions for future research.
Acknowledgement
I would like to express my deepest gratitude to my supervisor, Hiroki Matui, for his continuous guidance and encouragement. I am also grateful to Kota Takeuchi and all members of his laboratory for their helpful comments and discussions through seminars. Finally, I would like to thank Tatsuya Goto, who introduced this problem at the Model Theory Summer School and Workshop 2024.
2 Preliminaries
In this section, we recall the model-theoretic framework for operator algebras.
We denote ordinals by $\alpha,\beta,\gamma,\dots$ and cardinals by $\kappa,\lambda,\dots$. For a cardinal $\kappa$, we denote by $\kappa^{+}$ the successor cardinal of $\kappa$, that is, the least cardinal greater than $\kappa$.
We denote the order type of the $\alpha$-th aleph number $\aleph_{\alpha}$ by $\omega_{\alpha}$. In particular, $\omega_{0}$ — the order type of $\aleph_{0}$, which is the set of all natural numbers — is simply denoted by $\omega$.
[…]
2.1 Languages and Structures
Definition 2.1.
A metric language $\mathcal{L}=\langle{\mathcal{S},\mathcal{C},\mathcal{F},\mathcal{R}}\rangle$ is given by specifying the following data:
  • A set $\mathcal{S}$ whose elements are called sorts.
  • A set of function symbols $\mathcal{F}$. For each function symbol $f\in\mathcal{F}$, there are an assigned (finite) tuple of sorts $\operatorname{dom}(f)=(S_{1},\dots,S_{n})$ called the domain of $f$ and an assigned sort $S_{f}=S\in\mathcal{S}$ called the co-domain of $f$. We write $f\colon S_{1}\times S_{2}\times\dots\times S_{n}\to S$.
  • For each function symbol $f\in\mathcal{F}$, there is an assigned function $\delta^{f}\colon(0,1]\to(0,1]$ called the moduli of uniform continuity of $f$.
  • A set of relation (or predicate) symbols $\mathcal{R}$. For each relation symbol $R\in\mathcal{R}$, there is an assigned tuple of sorts $\operatorname{dom}(R)=(S_{1},\dots,S_{n})$, a closed interval $\operatorname{cod}(R)=I_{R}\subseteq\mathbb{R}$ and a function $\delta^{R}\colon(0,1]\to(0,1]$. We write $R\colon S_{1}\times S_{2}\times\dots\times S_{n}\to I_{R}$.
  • A set of constant symbols $\mathcal{C}$. For each constant symbol $c\in\mathcal{C}$, there is an assigned sort $S_{c}\in\mathcal{S}$. One can think of constants as 0-ary function symbols.
  • For each sort $S\in\mathcal{S}$, there is a relation symbol $d_{S}\colon S\times S\to[0,B_{S}]$ intended to be a metric, and a distinct (sufficiently large) infinite set whose elements are called variable symbols with respect to $S$.
Definition 2.2.
Let $\mathcal{L}=\langle{\mathcal{S},\mathcal{C},\mathcal{F},\mathcal{R}}\rangle$ be a metric language. An $\mathcal{L}$-structure $M$ is given by the following data:
  • For each sort $S\in\mathcal{S}$, $(S^{M},d_{S}^{M})$ is a complete (non-empty) metric space of diameter is not greater than $B_{S}$. We will write $\bar{a}\in M$ to mean $\bar{a}=(a_{1},a_{2},\dots,a_{n})\in\prod_{i}S_{i}^{M}$ for some sorts $S_{i}$ of $\mathcal{L}$.
  • For each function symbol $f\colon\prod_{i}S_{i}\to S\in\mathcal{F}$, its interpretation is given by a uniformly continuous function $f^{M}\colon\prod_{i}S_{i}^{M}\to S^{M}$, with $\delta^{f}$ as its modulus of uniform continuity. That is, for all $\varepsilon>0$ and $\bar{a},\bar{b}\in M$ with $\max_{1\leq i\leq n}d_{S_{i}}^{M}(a_{i},b_{i})<\delta^{f}(\varepsilon)$, we have $d_{S}^{M}(f^{M}(\bar{a}),f^{M}(\bar{b}))\leq\varepsilon$.
  • For each relation symbol $R\colon\prod_{i}S_{i}\to I_{R}\in\mathcal{R}$, its interpretation is given by a uniformly continuous real-valued function $R^{M}\colon\prod_{i}S_{i}^{M}\to I_{R}$, with $\delta^{R}$ as its modulus of uniform continuity. That is, for all $\varepsilon>0$ and $\bar{a},\bar{b}\in M$ with $\max_{1\leq i\leq n}d_{S_{i}}^{M}(a_{i},b_{i})<\delta^{f}(\varepsilon)$, we have $\left\lvert{R^{M}(\bar{a})-R^{M}(\bar{b})}\right\rvert\leq\varepsilon$.
  • For each constant symbol $c\in\mathcal{C}$, its interpretation $c^{M}$ is an element of $S_{c}^{M}$.
We note that the metric function $d$ of a metric space $(X,d)$ is uniformly continuous with $\delta(\varepsilon)=\varepsilon/2$ as a modulus of uniform continuity for $d$.
A $\mathrm{C}^{\ast}$}-algebra is a complex Banach algebra $A$ equipped with an involution $\ast\colon A\to A$ (i.e., a map satisfying $(\lambda a+b)^{\ast}=\bar{\lambda}a^{\ast}+b^{\ast}$, $(a^{\ast})^{\ast}=a$, and $(ab)^{\ast}=b^{\ast}a^{\ast}$ for all $a,b\in A$ and $\lambda\in\mathbb{C}$) such that the $\mathrm{C}^{\ast}$}-identity $\left\lVert{a^{\ast}a}\right\rVert=\left\lVert{a}\right\rVert^{2}$ holds for all $x\in A$. $\mathrm{C}^{\ast}$}-algebras are not bounded, so they cannot be regarded as metric structures as they are. To address this, we introduce the following language and structure so that they can be treated as metric structures.
Example 2.3 (unital $\mathrm{C}^{\ast}$}-algebras).
The language of unital $\mathrm{C}^{\ast}$}-algebras $\mathcal{L}_{{\mathrm{C}^{\ast}}}$ is given by the following data:
  • The set of sorts of $\mathcal{L}_{{\mathrm{C}^{\ast}}}$ is $\mathcal{S}_{{\mathrm{C}^{\ast}}}=\left\{S_{n}:n<\omega\right\}$ where $B_{S_{n}}=2n$ for each $n<\omega$.
  • $\mathcal{F}_{{\mathrm{C}^{\ast}}}$ consists of the following symbols:
    • $\circ$
      $_{n}\colon S_{n}\times S_{n}\to S_{2n}$ for each $n<\omega$,
    • $\circ$
      $\cdot_{n}\colon S_{n}\times S_{n}\to S_{n^{2}}$ for each $n<\omega$,
    • $\circ$
      $\lambda(\cdot)_{n}\colon S_{n}\to S_{\lceil\lambda\rceil n}$ for each $\lambda\in\mathbb{C}$ and $n<\omega$,
    • $\circ$
      $(\cdot)^{\ast}_{n}\colon S_{n}\to S_{n}$ for each $n<\omega$,
    • $\circ$
      $\iota_{n,m}\colon S_{n}\to S_{m}$ for each $n<m<\omega$.
  • $\mathcal{C}_{{\mathrm{C}^{\ast}}}$ consists of the symbols 0 and $1$ which belong to $S_{0}$ and $S_{1}$ respectively.
  • $\mathcal{R}_{{\mathrm{C}^{\ast}}}$ consists of the symbols $\left\lVert{\cdot}\right\rVert_{n}\colon S_{n}\to[0,n]$ for each $n<\omega$.
Let $A$ be a unital $\mathrm{C}^{\ast}$}-algebra. The $\mathcal{L}_{{\mathrm{C}^{\ast}}}$-structure $\mathrm{D}(A)$, called the dissection of $A$, is defined as follows:
  • For each $n<\omega$, $S_{n}^{\mathrm{D}(A)}$ is the closed unit ball of radius $n$.
  • Each interpretation of a function or relation symbol is a restriction of operations of $A$ for each sort.
  • $0^{\mathrm{D}(A)}$ and $1^{\mathrm{D}(A)}$ are the zero element and unit element of $A$.
2.2 Terms, Formulas and Their Interpretations
Hereafter, we let $\mathcal{L}$ denote a metric language. Terms and formulas of $\mathcal{L}$ are inductively defined.
Definition 2.4.
$\mathcal{L}$-terms along with their co-domain and variables are defined inductively as follows:
  1. (1)
    Each constant symbols $c\in\mathcal{C}$ are terms with no variables.
  2. (2)
    Each variable symbols $x$ are also terms.
  3. (3)
    If $f\colon\prod_{i}S_{i}\to S\in\mathcal{F}$ and $\tau_{1},\dots,\tau_{n}$ are terms with $\operatorname{cod}(\tau_{i})=S_{i}$, then the composition $\tau=f(\tau_{1},\dots,\tau_{n})$ is a term with $\operatorname{cod}(\tau)=\operatorname{cod}(f)$ and the variables of $\tau$ are the union of variables of $\tau_{i}$’s.
We often write a term as $\tau(\bar{x})$ to indicate that its variables are contained in the tuple of variables $\bar{x}$.
Definition 2.5.
$\mathcal{L}$-formulas along with their free variables are defined inductively as follows:
  1. (1)
    (Atomic) For a relation symbol $R\colon\prod_{i}S_{i}\to I_{R}\in\mathcal{R}$ and terms $\tau_{1},\dots,\tau_{n}$ with $\operatorname{cod}(\tau_{i})=S_{i}$, $R(\tau_{1},\dots,\tau_{n})$ is an $\mathcal{L}$-formula whose free variables are the union of variables of $\tau_{i}$’s. Formulas of this form are is called atomic formulas.
  2. (2)
    (Logical Conjunction) For $\mathcal{L}$-formulas $\varphi_{1},\dots,\varphi_{n}$ and a continuous function $g\colon\mathbb{R}^{n}\to\mathbb{R}$, the composition $g(\varphi_{1},\dots,\varphi_{n})$ is an $\mathcal{L}$-formula whose free variables are the union of free variables of $\varphi_{i}$’s.
  3. (3)
    (Quantification) For an $\mathcal{L}$-formula $\varphi$ and a variable symbol $x^{S}$, $\sup_{x\in S}\varphi$ and $\inf_{x\in S}\varphi$ are also $\mathcal{L}$-formulas whose free variables are those of $\varphi$ except $x$.
We often write a formula as $\varphi(\bar{x})$ to indicate that its free variables are contained in the tuple of variables $\bar{x}$. We note that every formula has only finitely many free variables.
We denote the set of all $\mathcal{L}$-formulas with $\bar{x}$ as the free variables by $\operatorname{Form}_{\bar{x}}(\mathcal{L})$, and set $\operatorname{Form}(\mathcal{L})\mathrel{:=}\bigcup_{\bar{x}}\operatorname{Form}_{\bar{x}}(\mathcal{L})$. An $\mathcal{L}$-formula with no free variables is called an $\mathcal{L}$-sentence. We denote the set of all $\mathcal{L}$-sentences by $\operatorname{Sent}(\mathcal{L})$.
Using the interpretation of symbols, we can define the interpretation of terms and formulas inductively.
Definition 2.6.
Let $M$ be an $\mathcal{L}$-structure. Then, for each $\mathcal{L}$-term $\tau=\tau(x_{1},\dots,x_{n})$, the interpretation of $\tau$, denoted by $\tau^{M}$, which is a function from $S_{1}^{M}\times\dots\times S_{n}^{M}$ to $S^{M}$ where $S_{i}$ is the sort of $x_{i}$ and $S$ is the co-domain of $\tau$, is defined inductively as follows:
  1. (1)
    If $\tau$ is a constant symbol $c\in\mathcal{C}$, then $\tau^{M}$ is $c^{M}$.
  2. (2)
    If $\tau$ is a variable symbols $x^{S}$, $\tau^{M}=x^{M}$ is defined by the identity map on $S$.
  3. (3)
    If $\tau=f(\tau_{1},\dots,\tau_{n})$, then $(f(\tau_{1},\dots,\tau_{n}))^{M}\mathrel{:=}f^{M}(\tau_{1}^{M},\dots,\tau_{n}^{M})$.
Similarly, for each $\mathcal{L}$-formula $\varphi=\varphi(x_{1},\dots,x_{n})$, the interpretation of $\varphi$, denoted by $\varphi^{M}$, which is a function from $S_{1}^{M}\times\dots\times S_{n}^{M}$ to $\mathbb{R}$ where $S_{i}$ is the sort of $x_{i}$, is defined inductively as follows:
  1. (1)
    If $\varphi=R(\tau_{1},\dots,\tau_{n})$, then $(R(\tau_{1},\dots,\tau_{n}))^{M}\mathrel{:=}R^{M}(\tau_{1}^{M},\dots,\tau_{n}^{M})$.
  2. (2)
    If $\varphi=g(\varphi_{1},\dots,\varphi_{n})$, then $(g(\varphi_{1},\dots,\varphi_{n}))^{M}\mathrel{:=}g^{M}(\varphi_{1}^{M},\dots,\varphi_{n}^{M})$.
  3. (3)
    If $\varphi=\sup_{x\in S}\psi$, then $(\sup_{x\in S}\psi)^{M}\mathrel{:=}\sup_{a\in S^{M}}\psi^{M}(a)$.
In particular, note that the interpretation of a sentence is given by a real number.
By definition, the interpretation of each $\mathcal{L}$-term is uniformly continuous, since $\mathcal{L}$-terms are built by composing function symbols, each of which is interpreted as uniformly continuous. Moreover, one can choose a modulus of uniform continuity that is independent of any $\mathcal{L}$-structure.
Similarly, the interpretation of each $\mathcal{L}$-formula is a uniformly continuous function, with a modulus of uniform continuity that is independent of any $\mathcal{L}$-structures.
Proposition 2.7.
For every $\mathcal{L}$-term $\tau(\bar{x})$, there is a function $\delta^{\tau}\colon(0,1]\to(0,1]$ which witnesses uniform continuity of $\tau^{M}$ for any $\mathcal{L}$-structure $M$. Similarly, for every $\mathcal{L}$-formula $\varphi(\bar{x})$, there exist a function $\delta^{\varphi}\colon(0,1]\to(0,1]$ and a bounded interval $I_{\varphi}\subseteq\mathbb{R}$ such that, for any $\mathcal{L}$-structure $M$, $\delta^{\varphi}$ witnesses uniform continuity of $\varphi^{M}$ and the image of $\varphi^{M}$ lies in $I_{\varphi}$ .
For details of the proof of section 2.2, see [MTFMS, Theorem 3.5].
2.3 Theories, Elementary Equivalences and Elementary Embeddings
Some (classical) model-theoretic concepts are generalized naturally to the continuous setting.
Definition 2.8.
Let $M$ be an $\mathcal{L}$-structure. For a subset $\mathrm{T}\subseteq\operatorname{Sent}(\mathcal{L})$, we say that $M$ is a model of $\mathrm{T}$, denoted by $M\models\mathrm{T}$, if $\varphi^{M}=0$ for all $\varphi\in\mathrm{T}$. When $\mathrm{T}$ is a singleton set $\mathrm{T}=\{\varphi\}$, we denote $M\models\{\varphi\}$ by $M\models\varphi$.
A set of $\mathcal{L}$-sentences that has a model is called an $\mathcal{L}$-theory. For an $\mathcal{L}$-structure $M$, $\operatorname{Th}(M)\mathrel{:=}\left\{\varphi\in\operatorname{Sent}(\mathcal{L}):M\models\varphi\right\}$ is called the theory of $M$.
Definition 2.9.
Let $M,N$ be $\mathcal{L}$-structures. We say that $M$ and $N$ are elementarily equivalent, denoted by $M\equiv N$, if $\operatorname{Th}(M)=\operatorname{Th}(N)$ holds. An $\mathcal{L}$-theory $\mathrm{T}$ is called complete if any two models of $\mathrm{T}$ are elementarily equivalent.
Example 2.10.
For any $\mathcal{L}$-structure $M$, the theory $\operatorname{Th}(M)$ is complete. Indeed, suppose that $N\models\operatorname{Th}(M)$ and $\varphi\in\operatorname{Th}(N)$. Since $\varphi^{M}-\varphi\in\operatorname{Th}(M)$, it follows that $0=(\varphi^{M}-\varphi)^{N}=\varphi^{M}-\varphi^{N}=\varphi^{M}$. Thus, we have $\varphi\in\operatorname{Th}(M)$.
Definition 2.11.
Let $M$ and $N$ be $\mathcal{L}$-structures. An $\mathcal{L}$-embedding $\rho\colon N\to M$ is a family of maps $(\rho_{S}\colon S^{N}\to S^{M})_{S\in\mathcal{S}}$ that satisfies the following conditions:
  • For each $c\in\mathcal{C}$, $\rho_{S_{c}}(c^{N})=c^{M}$ holds.
  • For each $f\colon\prod_{i}S_{i}\to S\in\mathcal{F}$ and $\bar{a}\in N$,
    $\rho_{S}(f^{N}(a_{1},a_{2},\dots,a_{n}))=f^{M}(\rho_{S_{1}}(a_{1}),\rho_{S_{2}}(a_{2}),\dots,\rho_{S_{n}}(a_{n}))$ (2.11.1)
    holds.
  • For each $R\colon\prod_{i}S_{i}\to I_{R}$ and $\bar{a}\in N$,
    $R^{N}(a_{1},a_{2},\dots,a_{n})=R^{N}(\rho_{S_{1}}(a_{1}),\rho_{S_{2}}(a_{2}),\dots,\rho_{S_{n}}(a_{n}))$ (2.11.2)
    holds.
If an embedding $\rho\colon N\to M$ is a family of inclusion maps, then we call that $N$ is a substructure of $M$.
We note that each $\rho_{S}$ is an isometric embedding between the metric spaces $(S^{N},d_{S}^{N})$ and $(S^{M},d_{S}^{M})$. In particular, each $\rho_{S}$ is injective.
Definition 2.12.
Let $M$ and $N$ be $\mathcal{L}$-structures. An embedding $\rho\colon N\to M$ is called an isomorphism if each $\rho_{S}$ is surjective. If there exists an isomorphism between $M$ and $N$, we say that $M$ and $N$ are isomorphic, and write $M\cong N$.
Note that if $\rho\colon N\to M$ is isomorphism, then the inverse $\rho^{-1}=(\rho_{S}^{-1})_{S\in\mathcal{S}}\colon M\to N$ is also isomorphism. Hence, the isomorphism relation is an equivalence relation.
By induction on the formation of terms and formulas, every embedding preserves the interpretation of all atomic formulas, and every isomorphism preserves the interpretation of all $\mathcal{L}$-formulas.
Definition 2.13.
Let $M,N$ be $\mathcal{L}$-structures. We say that an embedding $\rho\colon N\to M$ is elementary if, for every $\mathcal{L}$-formula $\varphi(\bar{x})$ and every $\bar{a}\in N$, we have
$\varphi^{N}(\bar{a})=\varphi^{M}(\rho(\bar{a})).$ (2.13.3)
If $N$ is a substructure of $M$ and the inclusion map is elementary, then we call $N$ an elementary substructure of $M$ and write $N\preceq M$.
As an example of an elementary embedding that is not an isomorphism, we will later see the diagonal embedding into an ultrapower (section 3.1).
Definition 2.14.
For an $\mathcal{L}$-theory $\mathrm{T}$, $\operatorname{Mod}(\mathrm{T})$ is the category with models of $\mathrm{T}$ as objects and elementary embeddings as morphisms. A class of $\mathcal{L}$-structures $\mathcal{C}$ is said to be elementary (or axiomatizable) if there exists an $\mathcal{L}$-theory $\mathrm{T}$ such that $\mathcal{C}=\operatorname{Mod}(\mathrm{T})$. We call such a theory $\mathrm{T}$ the theory of the class $\mathcal{C}$.
Example 2.15.
The class of all dissections of unital $\mathrm{C}^{\ast}$}-algebras (see section 2.1) is axiomatizable by explicitly writing out each condition. For a detailed list of these formulas, see [MToCA, Example 2.2.1]. We denote the theory of this class by $\mathrm{T}_{\mathrm{C}^{\ast}}$.
2.4 The Downward Löwenheim-Skolem-Tarski Theorem and Elementary chains
The downward Löwenheim-Skolem-Tarski theorem, a well-known result in the classical model theory for elementary substructures, generalizes naturally to the continuous setting.
Definition 2.16.
For a topological space $X$, we define density character of X, denoted by $\chi(X)$, as the smallest cardinality of a dense subset of $X$. We note that $X$ is called separable when $\chi(X)=\aleph_{0}$.
We also define the density character of a family of topological spaces $A=(S^{A})_{S\in\mathcal{S}}$, denoted by $\chi(A)$, to be $\sup_{S\in\mathcal{S}}\chi(S^{A})$.
Definition 2.17.
Let $\mathrm{T}$ be an $\mathcal{L}$-theory. We define the semi-norm $\left\lVert{\cdot}\right\rVert_{\mathrm{T}}$ on $\operatorname{Form}_{\bar{x}}(\mathcal{L})$ (for each $\bar{x}$) as follows:
$\left\lVert{\varphi}\right\rVert_{\mathrm{T}}\mathrel{:=}\sup\left\{\left\lvert{\varphi^{M}(\bar{a})}\right\rvert:M\models\mathrm{T},\bar{a}\in M\right\},\quad\varphi(\bar{x})\in\operatorname{Form}_{\bar{x}}(\mathcal{L}).$ (2.17.4)
We sometimes regard $\operatorname{Form}_{\bar{x}}(\mathcal{L})$ as the quotient space $\operatorname{Form}_{\bar{x}}(\mathcal{L})/\left\{\varphi:\left\lVert{\varphi}\right\rVert_{\mathrm{T}}=0\right\}$, and $\left\lVert{\cdot}\right\rVert_{\mathrm{T}}$ to be the norm on $\operatorname{Form}_{\bar{x}}(\mathcal{L})$.
For a set of $\mathcal{L}$-formulas $\Sigma(\bar{x})\subseteq\operatorname{Form}_{\bar{x}}(\mathcal{L})$, we denote by $\chi(\Sigma(\bar{x}),\mathrm{T})$ the density character of $\Sigma(\bar{x})$ with respect to the topology induced by the (semi-)norm $\left\lVert{\cdot}\right\rVert_{\mathrm{T}}$. We also denote by
$\chi(\mathcal{L},\mathrm{T})\mathrel{:=}\sum_{S_{1},\dots,S_{n}}\chi(\operatorname{Form}_{\bar{x}}(\mathcal{L}),\mathrm{T})$ (2.17.5)
where $S_{1},\dots,S_{n}$ run over all finitely many sorts in $\mathcal{L}$ and $\bar{x}=(x_{1},\dots,x_{n})$ is a finite tuple of variables of $S_{i}$’s. We say that $\mathcal{L}$ is separable relative to $\mathrm{T}$ if $\chi(\mathcal{L},\mathrm{T})\leq\aleph_{0}$.
Example 2.18.
By the Stone-Weierstrass theorem, every continuous function on a compact subset of $\mathbb{R}^{n}$ can be approximated by polynomials with rational coefficients. Thus, we have $\chi(\mathcal{L},\emptyset)\leq\left\lvert{\mathcal{L}}\right\rvert+\aleph_{0}$, where $\left\lvert{\mathcal{L}}\right\rvert$ denotes the sum of the cardinalities of the sets of symbols and sorts in $\mathcal{L}$.
Moreover, in the setting of unital $\mathrm{C}^{\ast}$}-algebras (section 2.3), it follows that $\mathcal{L}_{{\mathrm{C}^{\ast}}}$ is separable relative to $\mathrm{T}_{{\mathrm{C}^{\ast}}}$, since we may restrict scalar multiplication to $\mathbb{Q}+i\mathbb{Q}$.
For an $\mathcal{L}$-structure $M$, we write $A\subseteq M$ if $A$ is a family of closed subsets $A=(S^{A})_{S\in\mathcal{S}}$ of $M$, indexed by the sorts $\mathcal{S}$ of $\mathcal{L}$.
The following is the continuous model theoretic version of the downward Löwenheim-Skolem-Tarski theorem. A proof can be found in [MTFMS, 7.3 Proposition] for instance. Here, we give a proof for conmpleteness.
Theorem 2.19 (Downward Löwenheim-Skolem-Tarski).
Let $\mathrm{T}$ be an $\mathcal{L}$-theory, $M$ be a model of $\mathrm{T}$ and $A\subseteq M$. Then, there exists $N\preceq M$ such that $A\subseteq N$ and $\chi(N)\leq\chi(A)+\chi(\mathcal{L},\mathrm{T})$.
To prove this theorem, we use the continuous version of the Tarski-Vaught test, which is discussed in [MTFMS, 4.5 Proposition] for instance. Here, we give its statement and a proof in a slightly different way.
Proposition 2.20 (Tarski-Vaught test).
Let $M$ be an $\mathcal{L}$-structure and $N\subseteq M$. Then, the following are equivalent:
  1. (1)
    NN becomes a substructure of $M$ and $N\preceq M$ holds.
  2. (2)
    For any $\mathcal{L}$-formula $\varphi(\bar{x},y)$ and $\bar{a}\in N$,
    $\inf_{b\in S^{N}}\varphi^{M}(\bar{a},b)=\inf_{b\in S^{M}}\varphi^{M}(\bar{a},b)$ (2.20.6)
    holds.
Proof.
(1) $\implies$ (2) : For any $\mathcal{L}$-formula $\varphi(\bar{x},y)$, consider $\psi(\bar{x})\equiv\inf_{y}\varphi^{M}(\bar{x},y)$. Since $N\preceq M$, for any $\bar{a}\in N$ we have
$\inf_{b\in S^{N}}\varphi^{M}(\bar{a},b)=\inf_{b\in S^{N}}\varphi^{N}(\bar{a},b)=\psi^{N}(\bar{a})=\psi^{M}(\bar{a})=\inf_{b\in S^{M}}\varphi^{M}(\bar{a},b).$ (2.20.7)
(2) $\implies$ (1) : Firstly, we show that $N$ is closed under any interpretation of function symbols. Let $f\colon\prod_{i}S_{i}\to S$ be a function symbol of $\mathcal{L}$ and $\bar{a}\in N$. Consider the $\mathcal{L}$-formula $\varphi(\bar{x},y)\equiv d(f(\bar{x}),y)$, then we have
$\inf_{b\in S^{N}}d^{M}(f(\bar{a}),b)=\inf_{b\in S^{M}}d^{M}(f(\bar{a}),b)=0.$ (2.20.8)
Thus $d(f(\bar{a}),N)=0$, and so we get $f(\bar{a})\in N$ because $N$ is a closed subset of $M$. Since constant symbols can be regarded as 0-ary function symbols, it follows from the same argument used for function symbols that the interpretation of each constant symbol lies in $N$. Therefore, we conclude that $N$ can be considered a substructure of $M$.
Next, we show that $N\preceq M$. We can prove this statement by induction on the formation of $\mathcal{L}$-formulas. It suffices to consider only the case of $\varphi(\bar{x})\equiv\inf_{y}\psi(\bar{x},y)$. By the assumption of induction, $\psi^{N}(\bar{a},b)=\psi^{M}(\bar{a},b)$ holds for all $\bar{a},b\in N$. Hence we have
$\varphi^{N}(\bar{a})=\inf_{b\in S^{N}}\psi^{N}(\bar{a},b)=\inf_{b\in S^{N}}\psi^{M}(\bar{a},b)=\inf_{b\in S^{M}}\psi^{M}(\bar{a},b)=\varphi^{M}(\bar{a})$ (2.20.9)
for all $\bar{a}\in N$. Thus, we obtain $N\preceq M$. ∎
Remark 2.21.
We note that it suffices to verify condition (2) of section 2.4 on a dense subset of each space $\operatorname{Form}_{\bar{x},y}(\mathcal{L})$. Indeed, if $\varphi_{n}^{M}(\bar{x},y)$ converges uniformly to $\varphi^{M}(\bar{x},y)$, then $\inf_{b\in C}\varphi_{n}^{M}(\bar{a},b)$ converges uniformly to $\inf_{b\in C}\varphi^{M}(\bar{a},b)$ for any $C\subseteq S^{M}$. Therefore, if each $\varphi_{n}$ satisfies the condition (2), then $\varphi$ also satisfies condition (2).
Proof of section 2.4.
Let $\kappa=\chi(\mathcal{L},\mathrm{T})$ and $(\varphi_{\alpha}(\bar{x}_{\alpha},y))_{\alpha<\kappa}$ be a dense subset of $\operatorname{Form}(\mathcal{L})$. We construct an increasing sequence $(A_{n})_{n<\omega}$ of subsets of $M$ by induction, such that the following holds:
  • $A_{0}\mathrel{:=}A$,
  • for all $n<\omega$, $\alpha<\kappa$ and $\bar{a}\in A_{n}$, there exists $b\in A_{n+1}$ such that
    $\varphi_{\alpha}^{M}(\bar{a},b)<\inf_{c\in S^{M}}\varphi_{\alpha}^{M}(\bar{a},c)+1/n,$ (2.21.10)
  • $\chi(A_{n+1})\leq\chi(A_{n})+\chi(\mathcal{L},\mathrm{T})$ holds for all $n<\omega$.
Suppose that $A_{n}$ has been defined, and let $D_{n}$ be a dense subset of $A_{n}$ with $\left\lvert{D_{n}}\right\rvert=\chi(A_{n})$. For each $\alpha<\kappa$ and $\bar{a}\in D_{n}$, we can find $b=b_{\alpha,\bar{a}}\in M$ such that
$\varphi_{\alpha}^{M}(\bar{a},b)<\inf_{c\in S^{M}}\varphi_{\alpha}^{M}(\bar{a},c)+1/n.$ (2.21.11)
Let $A_{n+1}\mathrel{:=}A_{n}\cup\left\{b_{\alpha}:\alpha<\kappa,\bar{a}\in D_{n}\right\}$. Then, we have $\chi(A_{n+1})\leq\chi(A_{n})+\kappa$.
Let $N$ be the closure of $\bigcup_{n<\omega}A_{n}$. Then $N$ satisfies the Tarski-Vaught test for $(\varphi_{\alpha})_{\alpha<\kappa}$, and it follows that $N\preceq M$. Moreover, $\chi(N)\leq\chi(A)+\chi(\mathcal{L},\mathrm{T})+\aleph_{0}$ holds. ∎
By a similar argument to that used in the proof of the Tarski-Vaught test, we obtain the following result, which is analogous to the classical result concerning the union of elementary chains.
Definition 2.22.
Let $(I,<)$ be a linear ordered set and $(M_{i})_{i\in I}$ be a family of $\mathcal{L}$-structures. We say that $(M_{i})_{i\in I}$ is an elementary chain if $M_{i}\preceq M_{j}$ holds for all $i<j$ in $I$.
The following is stated in [MTFMS, 7.2 Proposition] with a sketchy proof. Here, we include a detailed proof for the reader’s convenience.
Proposition 2.23.
Suppose that $(M_{i})_{i\in I}$ is an elementary chain of $\mathcal{L}$-structures. Let $M$ be the completion of the union of $M_{i}$’s. Then $M_{i}\preceq M$ holds for all $i\in I$.
Proof.
It is easy to check that each interpretation of the symbols of $\mathcal{L}$ can be extended to the sorts of $M$. Thus, $M$ can be regarded as an $\mathcal{L}$-structure, with each $M_{i}$ being a substructure of $M$.
Let us prove $M_{i}\preceq M$ for all $i\in I$. We show $\varphi^{M_{i}}(\bar{a})=\varphi^{M}(\bar{a})$ for $\bar{a}\in M_{i}$ by induction on the formation of $\mathcal{L}$-formulas. Since $M_{i}$ is a substructure of $M$, this clearly holds for any atomic formulas. Also, formulas obtained by logical conjunctions can be easily handled.
Suppose that $\psi^{M_{i}}(\bar{a},b)=\psi^{M}(\bar{a},b)$ for every $i\in I$ and $\bar{a},b\in M_{i}$. Consider the formula $\varphi(\bar{x})\equiv\inf_{y}\psi(\bar{x},y)$. Let $\bar{a}\in M_{i}$ and $\varepsilon>0$. Then, there exists $b\in M$ such that $\psi^{M}(\bar{a},b)<\varphi^{M}(\bar{a})+\varepsilon/2$ holds. Furthermore, by the continuity of $\psi^{M}(\bar{x},y)$ and the density of $\bigcup_{i\in I}M_{i}$, we can take some $j\in I$ and $c\in M_{j}$ such that $i\leq j$ and $\psi^{M}(\bar{a},c)<\psi^{M}(\bar{a},b)+\varepsilon/2$ holds. Then, since $M_{i}\preceq M_{j}$, we have
$\varphi^{M_{i}}(\bar{a})=\varphi^{M_{j}}(\bar{a})\leq\psi^{M_{j}}(\bar{a},c)=\psi^{M}(\bar{a},c)<\varphi^{M}(\bar{a})+\varepsilon.$ (2.23.12)
Since $\varepsilon>0$ is arbitrary, we obtain $\varphi^{M_{i}}(\bar{a})\leq\varphi^{M}(\bar{a})$. The reverse inequality follows immediately from $M_{i}\subseteq M$. Hence, we have $\varphi^{M_{i}}(\bar{a})=\varphi^{M}(\bar{a})$. ∎
Remark 2.24.
If $(I,<)$ is an ordinal $\alpha$ with $\mathrm{cf}(\alpha)\neq\omega$ (in particular, when $\alpha$ is a regular uncountable cardinal such as $\omega_{1}$ and $\omega_{2}$), then the union of an elementary chain $(M_{\beta})_{\beta<\alpha}$ is automatically complete, since any Cauchy sequences are contained in some $M_{\beta}$.
3 Ultraproducts and Related Concepts
In this section, we introduce ultraproducts for metric structures and related concepts to ultraproducts.
3.1 Construction of the Ultraproducts
Firstly, we introduce construction of the ultraproducts for metric structures.
Before starting the discussion, we introduce ultrafilters and ultralimits.
Definition 3.1.
Let $I$ be a non-empty set. A (proper) filter on $I$ is a subset $\mathcal{F}\subseteq\mathcal{P}(I)$, where $\mathcal{P}(I)$ is the power set of $I$, that satisfies the following conditions:
  1. (1)
    $\mathcal{F}\neq\emptyset,\mathcal{P}(I)$.
  2. (2)
    For all $\in\mathcal{F}, ABA\cap B\in\mathcal{F}$ holds.
  3. (3)
    For all $A\in\mathcal{F}$ and $B\subseteq I$, if $A\subseteq B$ then $B\in\mathcal{F}$ holds.
If a filter $\mathcal{U}$ on $I$ satisfies the following condition, then we call $\mathcal{U}$ an ultrafilter on $I$:
  1. (4)
    For all $A\subseteq I$, either $A\in\mathcal{U}$ or $I\setminus A\in\mathcal{U}$ holds.
Let $\beta I$ be the set of all ultrafilters on $I$. For each $i\in I$, $i{\uparrow}\mathrel{:=}\left\{A\subseteq I\mathrel{}\middle|\mathrel{}i\in A\right\}$ is an ultrafilter on $I$ called a principal ultrafilter. By this correspondence $I\ni i\mapsto i{\uparrow}\in\beta I$, we can consider $I$ as a subset of $\beta I$. An element of $\beta I\setminus I$ is called a non-principal (or free) ultrafilter on $I$.
By Zorn’s lemma, for any filter $\mathcal{F}$ on $I$, there exists an ultrafilter $\mathcal{U}\in\beta I$ such that $\mathcal{F}\subseteq\mathcal{U}$. If $I$ is an infinite set, then $\mathcal{F}_{\mathrm{fin}}\mathrel{:=}\left\{A\subseteq I:\left\lvert{I\setminus A}\right\rvert<\aleph_{0}\right\}$ is a filter on $I$ (called the Fréchet filter). Thus, there exists an ultrafilter $\mathcal{U}\in\beta I$ such that $\mathcal{F}_{\mathrm{fin}}\subseteq\mathcal{U}$. Since all ultrafilters $\mathcal{U}\supseteq\mathcal{F}_{\mathrm{fin}}$ are non-principal, we have $\beta I\setminus I\neq\emptyset$ unless $I$ is finite.
Definition 3.2.
Suppose that $I$ is a non-empty set and let $\mathcal{U}\in\beta I$ be an ultrafilter. Let $X$ be a topological space, and let $(x_{i})_{i\in I}$ be a family of points in $X$ indexed by $I$. We say that a point $x\in X$ is the $\mathcal{U}$-limit of $(x_{i})_{i\in I}$, denoted by $x=\lim_{i\to\mathcal{U}}x_{i}$, if for every neighborhood $N$ of $x$, the set $\left\{i\in I:x_{i}\in N\right\}$ belongs to $\mathcal{U}$. We note that if $X$ is a compact Hausdorff space, then the $\mathcal{U}$-limit always exists and is unique.
Remark 3.3.
Suppose $\mathcal{U}$ is a principal ultrafilter $i_{0}{\uparrow}$ for some $i_{0}\in I$. Then $x_{i_{0}}$ is a $\mathcal{U}$-limit of $(x_{i})_{i\in I}$. Thus, the nontrivial case arises when $\mathcal{U}$ is non-principal.
Now we return to the discussion about the construction of the ultraproducts. Let $(M_{i})_{i\in I}$ be a family of $\mathcal{L}$-structures indexed by $I$, and $\mathcal{U}\in\beta I$. We define the $\mathcal{L}$-structure $\bar{M}$ by the following way. Firstly, we define the psudo-metric $\bar{d}_{S}$ on $\bar{S}\mathrel{:=}\prod_{i\in I}S^{M_{i}}$, for each sort $S\in\mathcal{S}$:
$\bar{d}_{S}((a_{i})_{i},(b_{i})_{i})\mathrel{:=}\lim_{i\to\mathcal{U}}d_{S}^{M_{i}}(a_{i},b_{i}),\quad(a_{i})_{i},(b_{i})_{i}\in\bar{S}.$ (3.3.1)
We define $S^{\bar{M}}$ as the quotient space of $\bar{S}$, with respect to the relation $a\sim b$ if $\bar{d}_{S}(a,b)=0$. Then, $S^{\bar{M}}$ becomes a complete metric space by the following (well-defined) metric $d_{S}^{\bar{M}}$:
$d_{S}^{\bar{M}}([(a_{i})_{i}],[(b_{i})_{i}])\mathrel{:=}\bar{d}_{S}((a_{i})_{i},(b_{i})_{i}),\quad[(a_{i})_{i}],[(b_{i})_{i}]\in S^{\bar{M}}.$ (3.3.2)
Similarly, the interpretation of each symbol in $\mathcal{L}$ can be (well-)defined as follows:
  • $c^{\bar{M}}\mathrel{:=}\left[\left(c^{M_{i}}\right)_{i}\right]$, for $c\in\mathcal{C}$.
  • $f^{\bar{M}}(a^{1},a^{2},\dots,a^{n})\mathrel{:=}\left[\left(f^{M_{i}}(a^{1}_{i},a^{2}_{i},\dots,a^{n}_{i})\right)_{i}\right]$, for $f\in\mathcal{F}$ and $a^{k}=[(a^{k}_{i})_{i}]\in S_{k}^{\bar{M}}$.
  • $R^{\bar{M}}(a^{1},a^{2},\dots,a^{n})\mathrel{:=}\lim\limits_{i\to\mathcal{U}}R^{M_{i}}(a^{1}_{i},a^{2}_{i},\dots,a^{n}_{i})$, for $R\in\mathcal{R}$ and $a^{k}=[(a^{k}_{i})_{i}]\in S_{k}^{\bar{M}}$.
By the above construction, $\bar{M}$ becomes an $\mathcal{L}$-structure. That is, for each $f\in\mathcal{F}$ and $R\in\mathcal{R}$, $f^{\bar{M}}$ and $R^{\bar{M}}$ are uniformly continuous functions with respect to $\delta^{f}$ and $\delta^{R}$.
This $\mathcal{L}$-structure $\bar{M}$} is called the ultraproduct of $(M_{i})_{i\in I}$ (with respect to $\mathcal{U}$), denoted by $\prod_{i\to\mathcal{U}}M_{i}$. In particular, when for each $M_{i}$ are the same structure $M$, the $\bar{M}$ is called the ultrapower of $M$ (with respect to $\mathcal{U}$), and denoted by $M^{\mathcal{U}}$.
Remark 3.4.
The ultrapower of a $\mathrm{C}^{\ast}$}-algebra AA is typically defined as
$A^{\mathcal{U}}\mathrel{:=}\ell^{\infty}(I,A)/\left\{(a_{i})_{i}:\lim\nolimits_{i\to\mathcal{U}}\left\lVert{a_{i}}\right\rVert=0\right\}.$ (3.4.3)
Evidently, the dissection of the ultrapower of a $\mathrm{C}^{\ast}$}-algebra agrees with the ultrapower of the dissection of it, that is $D(A^{\mathcal{U}})\cong D(A)^{\mathcal{U}}$ (see [FHS2:2014, Proposition 4.1]).
One of the most important model-theoretic properties of ultraproducts is the following statement called Łoś’ theorem.
Theorem 3.5 (Łoś’ theorem).
Let $(M_{i})_{i\in I}$ be a family of $\mathcal{L}$-structures, $\mathcal{U}\in\beta I$, and $\bar{M}=\prod_{i\to\mathcal{U}}M_{i}$. Then, for every $\mathcal{L}$-formula $\varphi(x_{1},\dots,x_{n})$ and $a^{k}=\left[\left(a^{k}_{i}\right)_{i}\right]\in S_{k}^{\bar{M}}$, the following equation holds:
$\varphi^{\bar{M}}(a^{1},a^{2},\dots,a^{n})=\lim_{i\to\mathcal{U}}\varphi^{M_{i}}(a^{1}_{i},a^{2}_{i},\dots,a^{n}_{i}).$ (3.5.4)
The proof of this theorem is by a straightforward induction on the formation of $\mathcal{L}$-formulas (see [MTFMS, Theorem 5.4]).
Corollary 3.6.
Let $M$ be an $\mathcal{L}$-structure and $\mathcal{U}\in\beta I$. Then the diagonal embedding $\iota\colon M\to M^{\mathcal{U}},a\mapsto\left[(a)_{i}\right]$ becomes an elementary embedding. That is, $M\cong\iota(M)\preceq M^{\mathcal{U}}$ holds. ∎
Remark 3.7.
Suppose that $M$ is an $\mathcal{L}$-structure such that every $S^{M}\ (S\in\mathcal{S})$ is compact. Then, $M$ is isomorphic to any ultrapower $M^{\mathcal{U}}$ via the embedding given by section 3.1.
Indeed, let $a=[(a_{i})_{i}]\in S^{M^{\mathcal{U}}}$. By the compactness of $S^{M}$, there exists $b\mathrel{:=}\lim_{i\to\mathcal{U}}a_{i}$ Then we have $\lim_{i\to\mathcal{U}}d_{S}^{M}(b,a_{i})=0$ i.e. $a=\iota(b)$.
The remark above means that, what we are interested in are the non-compact cases. In the case of operator algebras, this corresponds to the infinite-dimensional algebras.
At the end of this section, we introduce one of the most important results concerning ultrapowers, the Keisler-Shelah theorem.
Theorem 3.8 (Keisler-Shelah theorem).
Suppose that $M$ and $N$ are $\mathcal{L}$-structures. Then, the following are equivalent :
  • $M\equiv N$.
  • There exists a set $I$ and $\mathcal{U},\mathcal{V}\in\beta I\setminus I$ such that $M^{\mathcal{U}}\cong N^{\mathcal{V}}$.
A proof of the Keisler-Shelah theorem in the continuous setting can be found in [GK:2022].
3.2 Compactness Theorem
In this subsection, we prove the compactness theorem by using the idea of the ultraproduct. The compactness theorem is related to the notion of types, which will be discussed in the next subsection.
Definition 3.9.
Let $\bar{x}=(x_{\alpha})_{\alpha<\kappa}$ be a (possibly infinite) family of variables in $\mathcal{L}$, and $\Sigma=\Sigma(\bar{x})$ be a collection of $\mathcal{L}$-formulas such that the free variables of each $\varphi\in\Sigma$ are contained in $\bar{x}$.
  • $\Sigma$ is said to be satisfiable if there exists an $\mathcal{L}$-structure $M$ and $\bar{a}=(a_{\alpha})_{\alpha<\kappa}\in M$ such that $\varphi^{M}(\bar{a})=0$ for all $\varphi(\bar{x})\in\Sigma$. In this case, we call $\Sigma$ is satisfiable in $M$.
  • $\Sigma$ is said to be finitely satisfiable if for every finite subset $\Sigma_{0}\subseteq\Sigma$ is satisfiable.
  • $\Sigma$ is said to be $\varepsilon$-approximately finitely satisfiable for $\varepsilon>0$ if for every finite subset $\Sigma_{0}\subseteq\Sigma$, there exists an $\mathcal{L}$-structure $M$ and $\bar{a}=(a_{\alpha})_{\alpha<\kappa}\in M$ such that $\left\lvert{\varphi^{M}(\bar{a})}\right\rvert<\delta$ for all $\varphi(\bar{x})\in\Sigma_{0}$.
Theorem 3.10 (compactness theorem).
Let $\bar{x}=(x_{\alpha})_{\alpha<\kappa}$ be a (possibly infinite) family of variables in $\mathcal{L}$, and $\Sigma=\Sigma(\bar{x})$ be a collection of $\mathcal{L}$-formulas with $\bar{x}$ as the free variables. Then, the following are equivalent:
  1. (1)
    $\Sigma$ is satisfiable.
  2. (2)
    $\Sigma$ is finitely satisfiable.
  3. (3)
    $\Sigma$ is $\varepsilon$-approximately finitely satisfiable for every $\varepsilon>0$.
Proof.
(1) $\implies$ (2) $\implies$ (3) is obvious. We show (3) $\implies$ (1).
Let $I\mathrel{:=}\left\{(\Sigma_{0},n)\mathrel{}\middle|\mathrel{}\Sigma_{0}\subseteq\Sigma:\text{finite},n<\omega\right\}$. By the assumption, for any $i=(\Sigma_{0},n)\in I$, there exists an $\mathcal{L}$-structure $M_{i}$ and $\bar{a}_{i}\in M_{i}$ such that $\left\lvert{\varphi(\bar{a}_{i})}\right\rvert<1/n$, for all $\varphi(\bar{x})\in\Sigma_{0}$.
For $\varphi\in\Sigma$ and $n<\omega$, we set $F_{\varphi,n}=\left\{i=(\Sigma_{0},m)\in I\mathrel{}\middle|\mathrel{}\varphi\in\Sigma_{0},n\leq m\right\}$. Then, the set $\mathcal{F}_{0}=\left\{F_{\varphi,n}\mathrel{}\middle|\mathrel{}\varphi\in\Sigma,n<\omega\right\}$ has the finite intersection property. Thus, there exists an ultrafilter $\mathcal{U}$ on $I$ which contains $\mathcal{F}_{0}$.
Consider $M\mathrel{:=}\prod_{i\to\mathcal{U}}M_{i}$. For every $\varphi\in\Sigma$ and $n<\omega$, since $F_{\varphi,n}\in\mathcal{U}$ and $\left\lvert{\varphi^{M_{i}}(\bar{a}_{i})}\right\rvert<1/n$ for $i\in F_{\varphi,n}$, we have $\lim\limits_{i\to\mathcal{U}}\varphi^{M_{i}}(\bar{a}_{i})=0$. Thus, by section 3.1, we obtain $\varphi^{M}([(\bar{a}_{i})_{i}])=0$, which implies $\Sigma$ is satisfiable in $M$. ∎
3.3 Types and Saturation
Definition 3.11.
Let $M$ be an $\mathcal{L}$-structure and $A\subseteq M$. The language $\mathcal{L}_{A}$ is the expansion of $\mathcal{L}$ obtained by adding a new constant symbol $c_{a}$ for each $a\in A$. The $\mathcal{L}_{A}$-structure $M_{A}$ is the expansion of $M$ by interpreting each $c_{a}$ as $a$ for all $a\in A$.
Remark 3.12.
We note that every $\mathcal{L}_{A}$-formula can be written in the form $\varphi(\bar{x},\bar{a})$, for some $\mathcal{L}$-formula $\varphi(\bar{x},\bar{y})$ and $\bar{a}\in A$. Thus, it follows that $\chi(\mathcal{L}_{A},\operatorname{Th}(M_{A}))\leq\chi(\mathcal{L},\operatorname{Th}(M))+\chi(A)+\aleph_{0}$
Definition 3.13.
Let $\mathrm{T}$ be an $\mathcal{L}$-theory and $\bar{x}$ be a finite tuple of variables in $\mathcal{L}$. A set of $\mathcal{L}$-formulas $t(\bar{x})$ with $\bar{x}$ as the free variables is called an ($\bar{x}$-)type of $\mathrm{T}$ if it is satisfiable in some model of $\mathrm{T}$. That is, there exists a model $M\models\mathrm{T}$ and a tuple $\bar{a}\in M$ such that $\varphi^{M}(\bar{a})=0$ holds for all $\varphi(\bar{x})$ in $t(\bar{x})$. Such a tuple $\bar{a}$ is called a realization of $t(\bar{x})$, and we say that $t(\bar{x})$ is realized by $\bar{a}$ in $M$.
Example 3.14.
Let $M$ be an $\mathcal{L}$-structure, $A\subseteq M$ and $\bar{b}\in M$. Then,
$\operatorname{tp}(\bar{b}/A)\mathrel{:=}\left\{\varphi(\bar{x},\bar{a})\in\operatorname{Form}_{\bar{x}}(\mathcal{L}_{A})\mathrel{}\middle|\mathrel{}\varphi^{M}(\bar{b},\bar{a})=0\right\}$ (3.14.5)
is a type of $\operatorname{Th}(M_{A})$. This $\operatorname{tp}(\bar{b}/A)$ is called the complete type of $\bar{b}$ over $A$.
A type of $\operatorname{Th}(M_{A})$ is usually referred to as a type over $A$. By the compactness theorem, types over $A\subseteq M$ are characterized by the following conditions:
Lemma 3.15.
Let $M$ be an $\mathcal{L}$-structure and $A\subseteq M$. For a set of $\mathcal{L}_{A}$-formulas $t(\bar{x})$ with $\bar{x}$ as the free variables, the following are equivalent :
  1. (1)
    $t(\bar{x})$ is a type over $A$.
  2. (2)
    $t(\bar{x})$ is $\varepsilon$-approximately finitely satisfiable in $M$ for every $\varepsilon>0$. That is, for all finite subset $t_{0}(\bar{x})\subseteq t(\bar{x})$ and $\varepsilon>0$, there exists $\bar{b}\in M$ such that $\left\lvert{\varphi^{M}(\bar{b})}\right\rvert<\varepsilon$ for all $\varphi(\bar{x})$ in $t_{0}(\bar{x})$.
Proof.
(2) $\implies$ (1) follows immediately from section 3.2. We show (1) $\implies$ (2).
By (1), there exists a model $N\models\operatorname{Th}(M_{A})$ such that $t(\bar{x})$ is realized in $N$. Fix a finite subset $t_{0}(\bar{x})\subseteq t(\bar{x})$. The $\mathcal{L}_{A}$-sentence
$\inf_{\bar{x}}\max\left\{\left\lvert{\varphi(\bar{x})}\right\rvert:\varphi(\bar{x})\in t_{0}(\bar{x})\right\}$ (3.15.6)
belongs to $\operatorname{Th}(N)$, and so it is in $\operatorname{Th}(M_{A})$ because $\operatorname{Th}(M_{A})$ is complete. Thus, for every $\varepsilon>0$, there exists some $\bar{b}\in M$ such that $\left\lvert{\varphi^{M}(\bar{b})}\right\rvert<\varepsilon$ holds for all $\varphi(\bar{x})$ in $t_{0}(\bar{x})$, which completes the proof of (2). ∎
We note that the compactness theorem places no restriction on the size of variables. Hence, by combining it with the downward Löwenheim-Skolem-Tarski theorem, we obtain the following result.
Proposition 3.16.
Suppose that $\mathrm{T}$ is an $\mathcal{L}$-theory and that $M$ is a model of $\mathrm{T}$. Let $(t_{i}(\bar{x}_{i}))_{i\in I}$ be a family of types over $M$. Then there exists $N\succeq M$ such that $\chi(N)\leq\chi(\mathcal{L},\mathrm{T})+\chi(M)+\left\lvert{I}\right\rvert$, and each $t_{i}(\bar{x}_{i})$ is realized in $N$.
Proof.
Suppose that $\chi(M)<\aleph_{0}$. Then $M$ is a finite metric space, and hence compact. By section 3.1, every type over $M$ is realized in $M$. Therefore, $N\mathrel{:=}M$ satisfies the condition of the statement.
Assume that $\chi(M)\geq\aleph_{0}$ and a family $(t_{i}(\bar{x}_{i}))_{i\in I}$ of types over $M$ is given. Let $\bar{x}\mathrel{:=}(\bar{x}_{i})_{i\in I}$ be a tuple of variables (Note that variable symbols should be chosen to avoid duplication), and define $\Sigma(\bar{x})\mathrel{:=}\bigcup_{i\in I}t_{i}(\bar{x}_{i})$.
For each finite subset $\Sigma_{0}(\bar{x})\subseteq\Sigma(\bar{x})$, by section 3.3, $\Sigma_{0}(\bar{x})$ is $\varepsilon$-approximately finitely satisfiable in $M$ for every $\varepsilon>0$. Thus, by the proof of section 3.2, $\Sigma(\bar{x})$ can be realized in some ultrapower of $M$. That is, there exists an elementary extension $N\succeq M$ and a tuple $\bar{a}\in N$ such that $\bar{a}$ realizes $\Sigma(\bar{x})$ in $N$.
Applying section 2.4 to the set $A\mathrel{:=}M\cup\left\{\bar{a}_{i}\mathrel{}\middle|\mathrel{}i\in I\right\}$ and using section 3.3, we may take $N$ such that $\chi(N)\leq\chi(\mathcal{L},T)+\chi(A)\leq\chi(\mathcal{L},\mathrm{T})+\chi(M)+\left\lvert{I}\right\rvert$. ∎
Definition 3.17.
Let $\kappa$ be a uncountable cardinal. We say that an $\mathcal{L}$-structure $M$ is $\kappa$-saturated if, for any $A\subseteq M$ such that $\chi(A)<\kappa$, every type over $A$ is realized in $M$.
An $\aleph_{1}$-saturated structure is also called countably saturated. A structure $M$ is said to be saturated if it is $\chi(M)$-saturated.
Remark 3.18.
The density character of any non-compact $\kappa$-saturated model $M$ is at least $\kappa$. Indeed, fix any subset $A\subseteq M$ with $\chi(A)<\kappa$. Since $M$ is not totally bounded, there exists $\varepsilon>0$ such that no finite collection of $\varepsilon$-balls covers $M$. Define the set $t(x)$ of $\mathcal{L}_{A}$-formulas by
$t(x)\mathrel{:=}\left\{\max(\varepsilon-d(x,a),0)\mathrel{}\middle|\mathrel{}a\in A\right\}.$ (3.18.7)
By the choice of $\varepsilon>0$, $t(x)$ is finitely satisfiable, and hence $t(x)$ is a type over $A$ by section 3.3. Since $M$ is $\kappa$-saturated, $t(x)$ is realized in $M$. Let $b\in M$ be a realization of $t(x)$. Then, since $d(b,a)\geq\varepsilon$ for all $a\in A$, we have $b\notin\overline{A}$. It follows that $A$ is not dense in $M$.
Lemma 3.19.
Let $\kappa$ be a uncountable cardinal and $M$ be a model of the $\mathcal{L}$-theory $\mathrm{T}$ with $\chi(\mathcal{L},\mathrm{T})<\kappa$. The following are equivalent:
  1. (1)
    $M$ is $\kappa$-saturated.
  2. (2)
    For any type $t(\bar{x})$ over $M$ with $\chi(t(\bar{x}),\mathrm{T})<\kappa$, $t(\bar{x})$ is realized in $M$.
Proof.
(1) $\implies$ (2) : Let $t(\bar{x})$ be a type over $M$ with $\chi(t(\bar{x}),\mathrm{T})<\kappa$. By taking a dense subset of $t(\bar{x})$, we may assume that $\left\lvert{t(\bar{x})}\right\rvert<\kappa$. Since every $\mathcal{L}_{M}$-formula contains only finitely many constant symbols, the cardinality of elements of $M$ that appear in $t(\bar{x})$ as constant symbols is also less than $\kappa$. Therefore, by the saturation of $M$, $t(\bar{x})$ is realized in $M$.
(2) $\implies$ (1) : Suppose $A\subseteq M$ with $\chi(A)<\kappa$, and let $t(\bar{x})$ be a type over $A$. Since $\chi(\mathcal{L},\mathrm{T})<\kappa$ and $\chi(A)<\kappa$, we can choose a dense subset $\Sigma(\bar{x})$ of $\operatorname{Form}_{\bar{x}}(\mathcal{L}_{A})$ such that $\left\lvert{\Sigma(\bar{x})}\right\rvert\leq\chi(\mathcal{L},\mathrm{T})+\chi(A)+\aleph_{0}<\kappa$. By the density of $\Sigma(\bar{x})$, for each $\varphi(\bar{x})\in t(\bar{x})$ and $n<\omega$, there exists $\psi_{n}=\psi_{n,\varphi}(\bar{x})\in\Sigma(\bar{x})$ such that $\left\lVert{\varphi-\psi_{n}}\right\rVert_{\mathrm{T}}<1/n$. Then, for any $N\models\operatorname{Th}(M_{A})$ and $\bar{a}\in N$, we have
$\varphi^{N}(\bar{a})=0\iff\left\lvert{\psi_{n}^{N}(\bar{a})}\right\rvert\leq 1/n\quad(n<\omega).$ (3.19.8)
This means that the realization of $\{\varphi(\bar{x})\}$ is equivalent to the realization of the set
$\left\{\max\left\{\left\lvert{\psi_{n}(\bar{x})}\right\rvert-1/n,0\right\}\mathrel{}\middle|\mathrel{}n<\omega\right\}.$ (3.19.9)
Therefore, the realization of $t(\bar{x})$ is equivalent to the realization of the set
$t^{\prime}(\bar{x})\mathrel{:=}\left\{\max\{\left\lvert{\psi_{n,\varphi}}\right\rvert-1/n,0\}\mathrel{}\middle|\mathrel{}\varphi\in t(\bar{x}),n<\omega\right\}.$ (3.19.10)
Since $\left\lvert{t^{\prime}(\bar{x})}\right\rvert\leq\left\lvert{\Sigma(\bar{x})}\right\rvert+\aleph_{0}<\kappa$, by assumption (2), $t^{\prime}(\bar{x})$ is realized in $M$. Hence, $t(\bar{x})$ is also realized in $M$. ∎
Theorem 3.20.
Suppose that $M$ is an $\mathcal{L}$-structure and $\kappa$ is an infinite cardinal. Then, there exists an elementary extension $N\succeq M$ such that $N$ is $\kappa^{+}$-saturated and $\chi(N)\leq 2^{\kappa}$.
Proof.
If necessary, by choosing $\kappa$ sufficiently large, we may assume that $\chi(\mathcal{L},\operatorname{Th}(M))<\kappa$. We inductively define an elementary chain $(M_{\alpha})_{\alpha<\kappa^{+}}$ such that:
  1. (1)
    $M_{0}=M$,
  2. (2)
    if $\gamma<\kappa^{+}$ is a limit ordinal, then $M_{\gamma}$ is the completion of $\bigcup_{\alpha<\gamma}M_{\alpha}$,
  3. (3)
    for every $\alpha<\kappa^{+}$, $M_{\alpha}\preceq M_{\alpha+1}$, $\chi(M_{\alpha})\leq 2^{\kappa}$,
  4. (4)
    for every type $t(\bar{x})$ over $M_{\alpha}$ with $\left\lvert{t(\bar{x})}\right\rvert\leq\kappa$, $t(\bar{x})$ is realized in $M_{\alpha+1}$.
Suppose that $\alpha<\kappa$, and let $N\mathrel{:=}M_{\alpha}$. Since $\chi(\mathcal{L},\operatorname{Th}(M))<\kappa$ and $\chi(N)\leq 2^{\kappa}$, it follows from section 3.3 that $\chi(\mathcal{L}_{N},\operatorname{Th}(N_{N}))\leq 2^{\kappa}$. Therefore, we can take a dense subset $\Sigma$ of $\operatorname{Form}(\mathcal{L}_{N})$ relative to $\operatorname{Th}(N_{N})$, with $\left\lvert{\Sigma(\bar{x})}\right\rvert\leq 2^{\kappa}$. To obtain $M_{\alpha+1}$, it is enough to consider all types $t(\bar{x})\subseteq\Sigma$ with $\left\lvert{t(\bar{x})}\right\rvert\leq\kappa$. Since $(2^{\kappa})^{\kappa}=2^{\kappa\times\kappa}=2^{\kappa}$, by using section 3.3 for these types, we can get $M_{\alpha+1}$ with $\chi(M_{\alpha+1})\leq\chi(N)+\chi(\mathcal{L}_{N},\operatorname{Th}(N_{N}))+2^{\kappa}=2^{\kappa}$.
Therefore, by section 2.4 and section 2.4, the union of the elementary chain $(M_{\alpha})_{\alpha<\kappa^{+}}$ is an elementary extension of $M=M_{0}$, and is $\kappa^{+}$-saturated due to the regularity of $\kappa^{+}$. ∎
We present some results concerning saturation and ultrapowers.
Proposition 3.21.
Let $M$ be an $\mathcal{L}$-structure with $\chi(\mathcal{L},\operatorname{Th}(M))\leq\aleph_{0}$ and $\mathcal{U}\in\beta\omega\setminus\omega$. Then, the ultrapower $M^{\mathcal{U}}$ is $\aleph_{1}$-saturated.
Proof.
By section 3.3, it suffices to consider a type $t(\bar{x})$ over $M$ such that $\left\lvert{t(\bar{x})}\right\rvert\leq\aleph_{0}$. Let $t(\bar{x})=\left\{\varphi_{n}(\bar{x}):n<\omega\right\}$. For each $n<\omega$, we can take $\bar{b}^{n}=[(\bar{b}_{i}^{n})_{i}]\in M^{\mathcal{U}}$ such that $\left\lvert{\varphi_{k}^{M^{\mathcal{U}}}(\bar{b}^{n})}\right\rvert<1/n$ for all $k<n$ by section 3.3. Thus, by section 3.1,
$Y_{n}\mathrel{:=}\left\{i\in\omega:\max_{k<n}\left\lvert{\varphi_{k}^{M}(\bar{b}_{i}^{n})}\right\rvert<1/n\right\}$ (3.21.11)
belongs to $\mathcal{U}$. For each $n<\omega$, we define $Z_{n}\mathrel{:=}\bigcap_{k=0}^{n}Y_{k}\setminus\{0,\dots,n\}$. Then, $(Z_{n})_{n<\omega}$ satisfies $Z_{n+1}\subseteq Z_{n}$ and $Z_{n}\in\mathcal{U}$ for all $n<\omega$. Moreover, we have $\bigcap_{n<\omega}Z_{n}=\emptyset$. Hence, for each $i<\omega$, there exists a unique $n(i)<\omega$ such that $i\in Z_{n(i)}\setminus Z_{n(i)+1}$. We then define $\bar{c}_{i}\mathrel{:=}\bar{b}^{n(i)}_{i}$ for $i<\omega$ and $\bar{c}\mathrel{:=}[(\bar{c}_{i})_{i}]$. We will show that $c$ realizes $t(x)$.
For each $k<n<\omega$, since $Z_{n}\subseteq Y_{n}$, we have $\left\lvert{\varphi_{k}^{M}(\bar{c}_{i})}\right\rvert<1/n$ for each $i\in Z_{n}$. Thus, for a fixed $k<\omega$ and for each $n>k$,
$X_{n}\mathrel{:=}\left\{i<\omega:\left\lvert{\varphi_{k}^{M}(\bar{c}_{i})}\right\rvert<1/n\right\}$ (3.21.12)
belongs to $\mathcal{U}$, Hence, by the definition of the $\mathcal{U}$-limit, we obtain $\lim_{i\to\mathcal{U}}\varphi_{k}^{M}(\bar{c}_{i})=0$. Therefore, $\bar{c}$ realizes $t(\bar{x})$. ∎
More generally, for every infinite cardinal $\kappa$, there is an ultrafilter $\mathcal{U}$ on $\kappa$ such that for any $\mathcal{L}$-structure $M$ with $\chi(\mathcal{L},\operatorname{Th}(M))\leq\kappa$, $M^{\mathcal{U}}$ is $\kappa^{+}$-saturated (see [Bradd:2023, Proposition 7.10]).
3.4 Definable Predicates and Definable Sets
The definable sets in the continuous setting are somewhat distinct from those in the classical setting.
Definition 3.22.
Let $\mathcal{L}$ be a metric language. A definable predicate (or definable formula) with respect to an $\mathcal{L}$-theory $\mathrm{T}$ is an element of the completion of (the quotient space of) $\operatorname{Form}_{\bar{x}}(\mathcal{L})$ with respect to the norm $\left\lVert{\cdot}\right\rVert_{\mathrm{T}}$. We denote the set of all definable predicates with respect to $\mathrm{T}$ by $\operatorname{Form}(\mathcal{L},\mathrm{T})$.
For definable predicate $\varphi(\bar{x})$ with respect to $\mathrm{T}$ and $\bar{a}\in M\models\mathrm{T}$, we define the interpretation of $\varphi(\bar{x})$ by $\varphi^{M}(\bar{a})\mathrel{:=}\lim_{n\to\infty}\varphi_{n}^{M}(\bar{a})$ where $(\varphi_{n}(\bar{x}))_{n}$ is a sequence of $\mathcal{L}$-formulas converging to $\varphi(\bar{x})$ with respect to $\left\lVert{\cdot}\right\rVert_{\mathrm{T}}$. Since the interpretation of $\mathcal{L}$-formulas is continuous and bounded, the interpretation $\varphi^{M}(\bar{x})$ of any definable predicate $\varphi(\bar{x})$ is also continuous and bounded. Moreover, definable predicates are closed under the taking logical conjunction and quantification (in the sense of section 2.2).
Several statements (e.g., section 3.1) can also be extended to the case of definable predicates.
Definition 3.23.
Let $\mathrm{T}$ be an $\mathcal{L}$-theory. A uniform assignment relative to the elementary class $\operatorname{Mod}(\mathrm{T})$ (or relative to the theory $\mathrm{T}$) is a functor $X$ from $\operatorname{Mod}(\mathrm{T})$ to the class of metric spaces that satisfies the following conditions:
  • There exist $S_{1},S_{2},\dots,S_{n}\in\mathcal{S}$ such that for every $M\models\mathrm{T}$, $X(M)$ is a closed subset of $\prod_{i}S_{i}^{M}$.
  • For each elementary embedding $\rho\colon N\to M$ in $\operatorname{Mod}(\mathrm{T})$, $X(\rho)$ is the restriction of $\rho$ to $X(N)$.
Such a functor $X$ is called a definable set if, there exists a definable predicate $\varphi(\bar{x})$ respect to $\mathrm{T}$ such that
$\varphi^{M}(\bar{a})=d^{M}(\bar{a},X(M))\mathrel{:=}\inf\left\{d^{M}(\bar{a},\bar{x})\mathrel{}\middle|\mathrel{}\bar{x}\in X(M)\right\}$ (3.23.13)
holds for all $M\models\mathrm{T}$ and $\bar{a}\in M$.
Theorem 3.24.
Suppose $\mathrm{T}$ is an $\mathcal{L}$-theory and $X$ is a uniform assignment relative to $\mathrm{T}$. The following are equivalent:
  1. (1)
    $X$ is a definable set relative to $\mathrm{T}$.
  2. (2)
    For any $\mathcal{L}$-formula $\varphi(\bar{x},\bar{y})$, there exists a definable predicate $\psi(\bar{y})$ with respect to $\mathrm{T}$ such that
    $\psi^{M}(\bar{a})=\inf\left\{\varphi^{M}(\bar{x},\bar{a}):\bar{x}\in X(M)\right\}$ (3.24.14)
    holds for all $M\models\mathrm{T}$ and $\bar{a}\in M$.
  3. (3)
    There exists an $\mathcal{L}$-formula $\varphi(\bar{x})$ such that:
    • For all $M\models\mathrm{T}$, $X(M)=\left\{\bar{a}\in M:\varphi^{M}(\bar{a})=0\right\}$.
    • For every $\varepsilon>0$, there exists $\delta>0$ such that, for all $M\models\mathrm{T}$ and $\bar{a}\in M$, if $\varphi^{M}(\bar{a})<\delta$ then $d^{M}(\bar{a},X(M))\leq\varepsilon$.
For the proof of this theorem, see [MToCA, Lemma 3.2.5] or [Bradd:2023, Theorem 8.2].
Example 3.25.
For an $\mathcal{L}$-theory $\mathrm{T}$ and a fixed $\mathcal{L}$-term $\tau(\bar{x})$, the image of $\tau^{M}$ for each $M\models\mathrm{T}$ is a definable set, since $\inf_{\bar{x}\in M}d(\bar{a},\tau^{M}(\bar{x}))$ is an $\mathcal{L}$-formula, and hence is a definable predicate with respect to $\mathrm{T}$. For instance, by letting $\tau(x)=x+x^{*}$ (resp. $\tau(x)=xx^{*}$) one can see that the set of all self-adjoint (resp. positive) elements is a definable set with respect to $\mathrm{T}_{\mathrm{C}^{\ast}}$.
Another example of a definable set is the set of all projections, which is definable relative to $\mathrm{T}_{\mathrm{C}^{\ast}}$. This follows from the fact that the predicate $\varphi\equiv\left\lVert{x^{*}-x}\right\rVert+\left\lVert{x^{2}-x}\right\rVert$ satisfies condition (3) of section 3.4 (see [Bradd:2023, Example 8.5]).
We describe the relationship between definable sets and ultraproducts.
Proposition 3.26.
Let $\mathrm{T}$ be an $\mathcal{L}$-theory, and $X$ be a definable set relative to $\mathrm{T}$. Then, for any family $(M_{i})_{i\in I}$ of models of $\mathrm{T}$ and $\mathcal{U}\in\beta I$, we have
$X(M)=\prod_{i\to\mathcal{U}}X(M_{i}),$ (3.26.15)
where $M\mathrel{:=}\prod_{i\to\mathcal{U}}M_{i}$.
Proof.
Since $X$ is a definable set, we can take a definable predicate $d_{X}(\bar{x})$ that expresses the distance from $\bar{x}$ to $X$ in any model of $\mathrm{T}$. Suppose that $\bar{a}=[(\bar{a}_{i})_{i}]\in\prod_{i\to\mathcal{U}}X(M_{i})$. Then, by section 3.1, we have
$d_{X}^{M}(\bar{a})=\lim_{i\to\mathcal{U}}d_{X}^{M_{i}}(\bar{a}_{i})=0.$ (3.26.16)
Hence, we obtain $\bar{a}\in X(M)$.
Suppose that $\bar{a}=[(\bar{a}_{i})_{i}]\in X(M)$. For a fixed $\varepsilon>0$, by section 3.1, the set
$J\mathrel{:=}\left\{i\in I:d_{X}^{M_{i}}(a_{i})<\varepsilon\right\}$ (3.26.17)
belongs to $\mathcal{U}$. Thus, for each $i\in J$, there exist $\bar{b}_{i}\in X(M_{i})$ with $d^{M_{i}}(\bar{a}_{i},\bar{b}_{i})<\varepsilon$. For $i\in I\setminus J$, choose an arbitrary $\bar{b}_{i}\in M_{i}$, and define $\bar{b}\mathrel{:=}[(\bar{b}_{i})_{i}]\in\prod_{i\to\mathcal{U}}X(M_{i})$. Then we have
$d^{M}\left(\bar{a},\prod\nolimits_{i\to\mathcal{U}}X(M_{i})\right)\leq d^{M}(\bar{a},\bar{b})=\lim_{i\to\mathcal{U}}d^{M_{i}}(\bar{a}_{i},\bar{b}_{i})\leq\varepsilon.$ (3.26.18)
Since $\varepsilon>0$ is arbitrary, we obtain $d^{M}(\bar{a},\prod_{i\to\mathcal{U}}X(M_{i}))=0$. Therefore, as $\prod_{i\to\mathcal{U}}X(M_{i})$ is closed in $M$, $\bar{a}$ is in $\prod_{i\to\mathcal{U}}X(M_{i})$. ∎
4 Some Results under Continuum Hypothesis
It is well known that countable ultrapowers exhibit good behavior under the assumption of the Continuum Hypothesis (CH). In this section, we present several results in this direction. For simplicity, we assume throughout this section that the language $\mathcal{L}$ is separable with respect to the theory $\mathrm{T}$.
Definition 4.1 (partial elementary embedding).
Let $M$ and $N$ be $\mathcal{L}$-structures. A partial embedding $\rho\colon M\to N$ is a family of partial map $\rho=(\rho_{S}\colon S^{A}\to S^{N})_{S\in\mathcal{S}}$ for some closed $A\subseteq M$, which preserves interpretations of all symbols in $\mathcal{L}$ applied to elements of $A$. If a partial embedding $\rho$ from $A\subset M$ to $N$ preserves interpretations of all $\mathcal{L}$-formulas on $A$, that is, $\varphi^{M}(\bar{a})=\varphi^{N}(\rho(\bar{a}))$ holds for all $\mathcal{L}$-formulas $\varphi(\bar{x})$ and $\bar{a}\in A$, then $\rho$ is said to be a partial elementary embedding.
Proposition 4.2.
Suppose that $M$ and $N$ are saturated models of $\mathrm{T}$ that are elementarily equivalent and have the same uncountable density character $\kappa$. Then $M$ and $N$ are isomorphic.
Moreover, for any partial elementary embedding $\Phi\colon M\to N$, $\Phi$ can be extended to an isomorphism $\Phi\colon M\to N$.
Proof.
We prove this by the back-and-forth method.
Fix dense subsets $\{a_{\alpha}\}_{\alpha<\kappa}$ of $M$ and $\{b_{\alpha}\}_{\alpha<\kappa}$ of $N$. For each $\alpha\leq\kappa$, we define closed subsets $A_{\alpha}\subseteq M$ and partial embeddings $\Phi_{\alpha}\colon M\to N$ inductively as follows:
  • For all $\alpha\leq\kappa$, $\Phi_{\alpha}$ is elementary and has domain $A_{\alpha}$.
  • $a_{\alpha}\in A_{\alpha+1},b_{\alpha}\in\Phi_{\alpha+1}(A_{\alpha+1})$ holds for all $\alpha<\kappa$.
  • If $\beta<\alpha$, then $A_{\beta}\subseteq A_{\alpha}$ and $\Phi_{\alpha}$ is an extension of $\Phi_{\beta}$.
  • $\chi(A_{\alpha})\leq\kappa$ holds for all $\alpha<\kappa$.
Let $A_{0}\mathrel{:=}\emptyset$. Suppose that $A_{\alpha}$ has already been defined. Then, let $t(x)\mathrel{:=}\operatorname{tp}(a_{\alpha}/A_{\alpha})$. Since $\chi(A_{\alpha})<\kappa$ and $N$ is saturated, there exists a realization $b\in N$ of the type $t(x)$. Extend $\Phi_{\alpha}$ to $\Psi\colon A_{\alpha}\cup\{a_{\alpha}\}\to N$ by setting $\Psi(a_{\alpha})\mathrel{:=}b$. (If $a_{\alpha}$ already belongs to $A_{\alpha}$, this step may be skipped.) Clearly, we have $\Psi$ is elementary. Similarly, consider the type $t^{\prime}(x)\mathrel{:=}\operatorname{tp}(b_{\alpha}/\Psi(A_{\alpha}\cup\{a_{\alpha}\}))$. We then obtain a realization $a\in M$ of $t^{\prime}(x)$. Define $A_{\alpha+1}\mathrel{:=}A_{\alpha}\cup\{a_{\alpha},a\}$ and define $\Phi_{\alpha+1}\colon A_{\alpha+1}\to N$ as the extension of $\Psi$ by setting $\Phi_{\alpha+1}(a)\mathrel{:=}b_{\alpha}$. Clearly, one has $\Phi_{\alpha+1}$ is elementary. For a limit ordinal $\gamma\leq\kappa$, define $A_{\gamma}$ as the closure of $\bigcup_{\alpha<\gamma}A_{\alpha}$, and let $\Phi_{\gamma}$ be the isometric extension of $\bigcup_{\alpha<\gamma}\Phi_{\alpha}$. By section 2.2, $\varphi^{M}(\bar{x})$ and $\varphi^{N}(\bar{x})$ are continuous, and so we have $\Phi_{\gamma}$ is elementary. Then, by the construction of $A_{\alpha}$, we have $A_{\kappa}=M$ and $\Phi_{\kappa}(A_{\kappa})=N$, so $\Phi_{\kappa}$ is an isomorphism between $M$ and $N$.
In the case where a partial elementary embedding $\Phi\colon M\to N$ is given, the construction can be carried out by starting with $A_{0}$ as the domain of $\Phi$ and setting $\Phi_{0}=\Phi$. ∎
Corollary 4.3.
Let $M$ and $N$ be non-compact models of $\mathrm{T}$ that are elementarily equivalent and satisfy $\chi(M),\chi(N)\leq{\mathfrak{c}}$. Then, for all $\mathcal{U},\mathcal{V}\in\beta\omega\setminus\omega$, the ultrapowers $M^{\mathcal{U}}$ and $N^{\mathcal{V}}$ are isomorphic.
Proof.
By section 3.3, $M^{\mathcal{U}}$ is ${\mathfrak{c}}=\aleph_{1}$-saturated under CH holds. Thus, by virtue of section 4, it suffices to show that the density character of $M^{\mathcal{U}}$ is equal to $\mathfrak{c}$}. By the construction of the ultraproduct, we have $\chi(M^{\mathcal{U}})\leq\chi(M)^{\aleph_{0}}\leq{\mathfrak{c}}^{\aleph_{0}}={\mathfrak{c}}$. On the other hand, by section 3.1, we can regard $M$ as a substructure of $M^{\mathcal{U}}$, and hence as a closed subset of $M^{\mathcal{U}}$. Thus, if $M$ is non-compact, then $M^{\mathcal{U}}$ is also non-compact. Therefore, by section 3.3, we have that $\chi(M^{\mathcal{U}})\geq\aleph_{1}={\mathfrak{c}}$. Hence, we obtain $\chi(M^{\mathcal{U}})={\mathfrak{c}}$. ∎
Hereafter, we consider the cardinality of isomorphisms between $M^{\mathcal{U}}$ and $N^{\mathcal{V}}$.
Lemma 4.4.
Let $M$ and $N$ be saturated models of $\mathrm{T}$. Suppose that $A\preceq M$ with $\chi(A)<\chi(M)$ and $\Phi\colon A\to N$ is an elementary embedding. Then, there exist two distinct partial elementary embeddings $\Phi_{0},\Phi_{1}\colon M\to N$ which both extend $\Phi$.
Moreover, suppose $t(x)$ is a type over $A$ such that the set $B$ of all realizations of $t(x)$ in $M$ is not equals to $A$. Then, there exist two partial elementary embeddings $\Phi_{0},\Phi_{1}\colon M\to N$ which extend $\Phi$ and differ on $B$.
Proof.
The first case is settled by taking $t(x)=\{d(x,x)\}$ and $B=M$, so we now consider the second case. Fix $b\in B\setminus A$ and let $\varepsilon\mathrel{:=}d(b,A)$. Consider the set of $\mathcal{L}_{A}$-formulas $t(x_{0},x_{1})$ defined by
$t(x_{0},x_{1})\mathrel{:=}\left\{\varphi(x_{i})\mathrel{}\middle|\mathrel{}\varphi(x)\in\operatorname{tp}(b/A),i=0,1\right\}\cup\{\max(\varepsilon-d(x_{0},x_{1}),0)\}$ (4.4.1)
Then $t(x_{0},x_{1})$ is a type over $A$. Indeed, for any finite subset $t^{\prime}(x_{0},x_{1})\subseteq t(x_{0},x_{1})$ and any $\delta>0$, write
$t_{0}(x_{0})\cup t_{1}(x_{1})=t^{\prime}(x_{0},x_{1})\setminus\{\max(\varepsilon-d(x_{0},x_{1}),0)\}.$ (4.4.2)
Since $b$ realizes $t_{1}(x_{1})$, the formula
$\psi\equiv\inf_{x_{1}}\min_{\varphi\in t_{1}}\left\lvert{\varphi(x_{1})}\right\rvert$ (4.4.3)
satisfies $\psi^{M}=0$. By $A\preceq M$, we also have $\psi^{A}=0$. Hence there exists $c\in A$ such that $\left\lvert{\varphi^{M}(c)}\right\rvert<\delta$ for all $\varphi(x_{1})\in t_{1}(x_{1})$. Moreover, $d(b,c)\geq d(b,c)=\varepsilon$, so the pair $(b,c)$ is $\delta$-approximately realization of $t^{\prime}(x_{0},x_{1})$. Therefore, by section 3.3, $t(x_{0},x_{1})$ is a type over $A$.
By the saturation of $N$, there exists a realization $(b_{0},b_{1})$ of $t(x_{0},x_{1})$ in $N$. As a result, we have partial elementary embeddings $\Phi_{0},\Phi_{1}\colon M\to N$ given by $\Phi_{i}(b)=b_{i}$ for $i=0,1$. ∎
By applying section 4 at each step $\alpha<\kappa$ in the proof of section 4, we obtain the following result:
Proposition 4.5.
Let $M$ and $N$ be saturated models of $\mathrm{T}$. Suppose $A\preceq M$ with $\chi(A)<\chi(M)$, and let $\Phi\colon A\to N$ be an elementary embedding. Then, there exist at least $2^{\chi(M)}$ isomorphisms from $M$ to $N$ that extend $\Phi$.
Moreover, suppose that $t(x)$ is a type over $A$ such that the set $B$ of all realizations of $t(x)$ in $M$ satisfies $\chi(B)>\chi(A)$. Then, there exist at least $2^{\chi(B)}$ isomorphisms from $M$ to $N$ that extend $\Phi$ and differ on $B$.
By using section 4 we obtain the following results in a similar fashion to section 4:
Corollary 4.6.
Assume CH holds. Suppose $M$ and $N$ are non-compact separable models of $\mathrm{T}$, and let $\Phi\colon M\to N$ be an isomorphism. Then, for any $\mathcal{U},\mathcal{V}\in\beta\omega\setminus\omega$, there exist $2^{\mathfrak{c}}$ isomorphisms from $M^{\mathcal{U}}$ to $N^{\mathcal{V}}$ that extend $\Phi$.
Moreover, suppose that $t(x)$ is a type over $A$ such that the set $B$ of all realizations of $t(x)$ in $M^{\mathcal{U}}$ has the density character $\mathfrak{c}$}. Then, there exist at least $2^{{\mathfrak{c}}}$ isomorphisms from $M^{\mathcal{U}}$ to $N^{\mathcal{V}}$ that extend $\Phi$ and differ on $B$.
We conclude by considering the case of ultraproducts of $\mathrm{C}^{\ast}$}-algebras. For an inclusion of $\mathrm{C}^{\ast}$}-algebras $A\subseteq M$, the relative commutant $A^{\prime}\cap M$ is the $\mathrm{C}^{\ast}$}-subalgebra of $M$ defined by
$A^{\prime}\cap M\mathrel{:=}\left\{x\in M\mathrel{}\middle|\mathrel{}xa=ax\ \text{for all}\ a\in A\right\}.$ (4.6.4)
The relative commutant can also be described as the set of all realizations of the type $t_{\mathrm{rc}}(x)\mathrel{:=}\left\{d(xa,ax)\mathrel{}\middle|\mathrel{}a\in A\right\}$. It is known that the relative commutant of an ultraproduct satisfies the following result:
Proposition 4.7 ([FHS1:2013, Lemma 5.2]).
Let $A$ be a infinite-dimensional separable unital $\mathrm{C}^{\ast}$}-algebra and $\mathcal{U}\in\beta\omega\setminus\omega$. Then, the relative commutant $A^{\prime}\cap A^{\mathcal{U}}$ is infinite-dimensional and non-separable.
Consequently, we obtain the following result by applying section 4 to the type $t_{\mathrm{rc}}(x)$.
Proposition 4.8 ([Farah:2019, Proposition 16.7.1]).
Assume CH holds. Let $A,B$ be infinite-dimensional separable unital $\mathrm{C}^{\ast}$}-algebras and $\Phi\colon A\to B$ an isomorphism. Then, for any $\mathcal{U},\mathcal{V}\in\beta\omega\setminus\omega$, there exist at least $2^{\mathfrak{c}}$ isomorphisms from $A^{\prime}\cap A^{\mathcal{U}}$ to $B^{\prime}\cap B^{\mathcal{V}}$.
Proof.
By section 4, there exist at least $2^{\mathfrak{c}}$ isomorphisms from $A^{\mathcal{U}}$ to $B^{\mathcal{V}}$ that extend $\Phi$ and differ on $A^{\prime}\cap A^{\mathcal{U}}$. Since isomorphisms preserve the set of realizations of a type setwise, the restriction of each such isomorphism to $A^{\prime}\cap A^{\mathcal{U}}$ yields an isomorphism from $A^{\prime}\cap A^{\mathcal{U}}$ to $B^{\prime}\cap B^{\mathcal{V}}$. Hence, we obtain at least $2^{\mathfrak{c}}$ isomorphisms from $A^{\prime}\cap A^{\mathcal{U}}$ to $B^{\prime}\cap B^{\mathcal{V}}$. ∎
5 Main Result
For a $\mathrm{C}^{\ast}$}-algebra AA, $\mathcal{P}(A)$ denotes the set of all projections in $A$. For $\in\mathcal{P}(A), pp$ is said to be minimal if, for any $q\in\mathcal{P}(A)$ with $0\leq q\leq p$ one has $q=0$ or $q=p$.
We let $\mathcal{C}$ denote the class of all unital $\mathrm{C}^{\ast}$}-algebras AA such that $\mathcal{P}(A)$ is commutative with respect to the product in $A$ and $A$ has no minimal projections.
Lemma 5.1.
The class $\mathcal{C}$ is axiomatizable.
Proof.
By section 3.4, $\mathcal{P}$ is definable. It follows from section 3.4 that
$\sup_{p\in\mathcal{P}}\sup_{q\in\mathcal{P}}\left\lVert{pq-qp}\right\rVert$ (5.1.1)
is a definable predicate relative to $\mathrm{T}_{\mathrm{C}^{\ast}}$. Thus, the commutativity of projections can be described by the $\mathcal{L}_{{\mathrm{C}^{\ast}}}$-formulas.
Suppose that $A$ is a unital $\mathrm{C}^{\ast}$}-algebra with $\mathcal{P}(A)$ commutative. For $p,q\in\mathcal{P}(A)$, the condition $q\leq p$ is equivalent to $pq=q$, and thus $\left\lVert{pq-q}\right\rVert=0$. The product of projections is again a projection since $\mathcal{P}(A)$ is commutative. Hence, if $\left\lVert{pq-q}\right\rVert\neq 0$, then $\left\lVert{pq-q}\right\rVert=\left\lVert{(1-p)q}\right\rVert=1$, for $p,q\in\mathcal{P}(A)$. Similarly, for all $p,q\in\mathcal{P}(A)$ with $q\leq p$, if $\left\lVert{p-q}\right\rVert\neq 0$, then $\left\lVert{p-q}\right\rVert=\left\lVert{p-pq}\right\rVert=1$. Therefore, $A\in\mathcal{C}$ if and only if the interpretation of the following definable predicate
$\sup_{p\in\mathcal{P}}\sup_{q\in\mathcal{P}}\min\{1-\left\lVert{pq-q}\right\rVert,\left\lVert{q}\right\rVert,\left\lVert{p-q}\right\rVert\}$ (5.1.2)
is equal to 0. Hence $\mathcal{C}$ is axiomatizable. ∎
Example 5.2.
Let $X$ be the Cantor space $2^{\omega}$. The $\mathrm{C}^{\ast}$}-algebra $A=C(X)$ is an abelian unital $\mathrm{C}^{\ast}$}-algebra with no minimal projections. More generally, for any abelian unital $\mathrm{C}^{\ast}$}-algebra $A=C(X)$, where $X$ is a compact Hausdorff space, $A$ has no minimal projections if and only if the Boolean algebra $\operatorname{clop}(X)$ consisting of all clopen subsets of $X$ is atomless.
A noncommutative example is obtained by tensoring a unital noncommutative $\mathrm{C}^{\ast}$}-algebra with no non-trivial projections such as the Jiang-Su algebra $\mathcal{Z}$. Namely, for $X$ as above, $\mathcal{Z}\otimes C(X)$ is in $\mathcal{C}$.
Theorem 5.3.
[…]
Proof.
Thanks to section 3.3, we may assume $M$ is $\kappa$-saturated for large enough $\kappa$.
For $p,q\in\mathcal{P}(M)$, we define the relation $q\dashv p$ by
$q\dashv p\mathrel{\overset{\text{def}}{\iff}}pq\neq 0\land(1-p)q\neq 0\iff 1-\min\{\left\lVert{pq}\right\rVert,\left\lVert{(1-p)q}\right\rVert\}=0.$ (5.3.3)
We define $p_{\alpha},q_{\alpha}\in M$ and $M_{\alpha}\preceq M$ inductively for ordinals $\alpha\ (\leq\omega_{2})$. One can find a separable elementaly submodel $M_{0}$ of $M$ by section 2.4. Suppose that $M_{\alpha}$ is already defined. We take $p_{\alpha},q_{\alpha}\in\mathcal{P}(M)\setminus\{0,1\}$ such that the following conditions hold:
$\left\{\begin{array}[]{ll}q\dashv p_{\alpha}&\text{for all}\ q\in\mathcal{P}(M_{\alpha})\setminus\{0,1\},\\ q_{\alpha}\not\dashv p&\text{for all}\ p\in\mathcal{P}(M_{\alpha})\setminus\{0,1\}.\end{array}\right$. (5.3.4)
To show existence of $p_{\alpha}\in\mathcal{P}(M)$, it is enough to check that the set
$t(x)\mathrel{:=}\{\left\lVert{x-x^{\ast}}\right\rVert,\left\lVert{x^{2}-x}\right\rVert\}\cup\left\{1-\min\{\left\lVert{xq}\right\rVert,\left\lVert{(1-x)q}\right\rVert\}:q\in\mathcal{P}(M_{\alpha})\setminus\{0,1\}\right\}$ (5.3.5)
is a type over $\mathcal{P}(M_{\alpha})\subseteq M$ because $M$ is $\kappa$-saturated. Let $n<\omega$ and $q_{i}\in\mathcal{P}(M_{\alpha})\setminus\{0,1\}\ (i<n)$. For each $\sigma\in 2^{n}$, we define
$q_{\sigma}\mathrel{:=}\prod_{i<n}\left[(1-\sigma(i))q_{i}+\sigma(i)(1-q_{i})\right].$ (5.3.6)
If $q_{\sigma}\neq 0$ choose $p_{\sigma}\in\mathcal{P}(M)$ with $0\lneq p_{\sigma}\lneq q_{\sigma}$. Then, the sum $p\mathrel{:=}\sum_{\sigma}p_{\sigma}$ satisfies $p\neq 0,1$ and $q_{i}\dashv p$ for all $i<n$ since $q_{i}=\sum_{\sigma:\sigma(i)=0}q_{\sigma}$ and $1-q_{i}=\sum_{\sigma:\sigma(i)=1}q_{\sigma}$. Thus, $t(x)$ is finitely satisfiable in $M$. It follows from section 3.3 that $t(x)$ is a type over $\mathcal{P}(M_{\alpha})$.
Similarly, to take $q_{\alpha}\in\mathcal{P}(M)$, it is enough to show that the set
$t^{\prime}(x)\mathrel{:=}\{\left\lVert{x-x^{\ast}}\right\rVert,\left\lVert{x^{2}-x}\right\rVert\}\cup\left\{\min\{\left\lVert{px}\right\rVert,\left\lVert{(1-p)x}\right\rVert\}:p\in\mathcal{P}(M_{\alpha})\setminus\{0,1\}\right\}$ (5.3.7)
Let $n<\omega$ and $p_{i}\in\mathcal{P}(M_{\alpha})\setminus\{0,1\}\ (i<n)$. For each $\sigma\in 2^{n}$, we define
$p^{\prime}_{\sigma}\mathrel{:=}\prod_{i<n}\left[(1-\sigma(i))p_{i}+\sigma(i)(1-p_{i})\right].$ (5.3.8)
Since $1=\sum_{\sigma}p^{\prime}_{\sigma}$ there exist $\sigma^{\prime}\in 2^{n}$ such that $q\mathrel{:=}p^{\prime}_{\sigma^{\prime}}\neq 0,1$. Then, we have $q\not\dashv p_{i}$ for all $i<n$. Thus, $t^{\prime}(x)$ is finitely satisfiable in $M$. It follows from section 3.3 that $t^{\prime}(x)$ is a type over $\mathcal{P}(M_{\alpha})$.
By section 2.4, we take $M_{\alpha}\preceq M$ such that $M_{\alpha}\cup\{p_{\alpha},q_{\alpha}\}\subseteq M_{\alpha+1}$ and $\chi(M_{\alpha+1})\leq\chi(M_{\alpha})$. When $\gamma$ is limit ordinals, we define $M_{\gamma}\mathrel{:=}\overline{\bigcup_{\alpha<\gamma}M_{\alpha}}$.
Let $A\mathrel{:=}M_{\omega_{1}}$ and $B\mathrel{:=}M_{\omega_{2}}$. Then, $A,B\preceq M$ by section 2.4. Furthermore, $A$ and $B$ satisfy the following conditions respectively.
Claim 1.
Suppose that $\mathcal{U}\in\beta\omega\setminus\omega$. Then the following condition holds:
$\exists P\subseteq\mathcal{P}(A^{\mathcal{U}})\setminus\{0,1\}\mathrel{\text{s.t.}}\left\lvert{P}\right\rvert\leq\omega_{1}\land\left[\forall q\in\mathcal{P}(A^{\mathcal{U}})\setminus\{0,1\},\exists p\in P,q\dashv p\right].$ (5.3.9)
(Proof of Claim 1).
Let $P\mathrel{:=}\left\{[(p_{\alpha})_{n}]:\alpha<\omega_{1}\right\}$, $\left\lvert{P}\right\rvert\leq\omega_{1}$ is obviously. We show that $P$ satisfies the above condition. We note that $\mathcal{P}(\cdot)$ is commutative with any ultraproducts, since $\mathcal{P}(\cdot)$ is definable. Thus, let $q\in\mathcal{P}(A^{\mathcal{U}})\setminus\{0,1\}=\mathcal{P}(A)^{\mathcal{U}}\setminus\{0,1\}$, then we can take $q=[(q_{n})_{n}]$ where $q_{n}\in\mathcal{P}(A)\setminus\{0,1\}$ for $n<\omega$. By section 2.4, we have $A=\bigcup_{\alpha<\omega_{1}}M_{\alpha}$. Hence, for each $n<\omega$, we can take $\alpha_{n}<\omega_{1}$ such that $q_{n}\in\mathcal{P}(M_{\alpha_{n}})$. Let $\alpha\mathrel{:=}\sup_{n<\omega}\alpha_{n}$, then $\alpha<\omega_{1}$ because $\omega_{1}$ is regular. Thus, we put $p\mathrel{:=}[(p_{\alpha})_{n}]\in P$, then $q_{n}\dashv p_{\alpha_{n}}$ for all $n<\omega$. Therefore, by the section 3.1, we obtain $q\dashv p$. ∎
Claim 2.
Suppose that $\mathcal{V}\in\beta\omega\setminus\omega$. Then the following condition holds:
$\forall P\subseteq\mathcal{P}(B^{\mathcal{V}})\setminus\{0,1\},\left\lvert{P}\right\rvert\leq\omega_{1}\implies\left[\exists q\in\mathcal{P}(B^{\mathcal{V}})\setminus\{0,1\},\forall p\in P,q\not\dashv p\right].$ (5.3.10)
(Proof of Claim 2).
Let $P\subseteq\mathcal{P}(B^{\mathcal{V}})\setminus\{0,1\}$ with $\left\lvert{P}\right\rvert\leq\omega_{1}$ is given. Let $P=\left\{r^{\alpha}:\alpha<\omega_{1}\right\}$, and $r^{\alpha}=[(r^{\alpha}_{n})_{n}]$ where $r^{\alpha}_{n}\in\mathcal{P}(B)\setminus\{0,1\}\ (n<\omega)$ for each $\alpha<\omega_{1}$. By section 2.4, we have $B=\bigcup_{\beta<\omega_{2}}M_{\beta}$. Hence, for each $n<\omega$ and $\alpha<\omega_{1}$, we can take $\beta^{\alpha}_{n}<\omega_{2}$ such that $r^{\alpha}_{n}\in M_{\beta^{\alpha}_{n}}$. Let $\beta\mathrel{:=}\sup\left\{\beta^{\alpha}_{n}:n<\omega,\alpha<\omega_{1}\right\}$ ($\omega_{2}$, since $\omega_{2}$ is regular) and $q\mathrel{:=}[(q_{\beta})_{n}]$. Then, for any $\alpha<\omega_{1}$, $q_{\beta}\not\dashv r^{\alpha}_{n}$ for all $n<\omega$ and thus we obtain $q\not\dashv r^{\alpha}$, by section 3.1. ∎
From the above claims, it follows that $A^{\mathcal{U}}$ and $B^{\mathcal{V}}$ are not isomorphic for all $\mathcal{U},\mathcal{V}\in\beta\omega\setminus\omega$.
[…]
6 Conclusion
section 5 assumes commutativity for the set of all projections in a $\mathrm{C}^{\ast}$}-algebra, which is a quite strong condition. It is natural to ask whether the same result holds for more general $\mathrm{C}^{\ast}$}-algebras.
Question 6.1.
[…]
We remark that, in general, when $\mathrm{T}$ is a stable $\mathcal{L}$-theory, it is known that any ultrapowers of $\mathcal{L}$-structures with density character $\leq{\mathfrak{c}}$ are isomorphic in ZFC ([FHS2:2014, Theorem 5.6 (1)]). Therefore, this question is negated when $\mathrm{T}$ is stable. However, any theory of infinite-dimensional unital $\mathrm{C}^{\ast}$}-algebras is unstable (see [FHS1:2013, Lemma 5.3] and [FHS2:2014, Theorem 5.5]). So, [FHS2:2014, Theorem 5.6 (1)] cannot be applied to this question.
In the construction given in the proof of section 5, the density characters of $A$ and $B$ are controlled only up to $\leq\aleph_{1}$ and $\leq\aleph_{2}$, respectively. Therefore, the case where $A$ and $B$ are separable is still unsolved.
Question 6.2.
[…]
To approach this problem, we would like to investigate the situation of section 5.
Definition 6.3.
Suppose that $X$ is a topological space. We say that $X$ is totally disconnected if any connected components of $X$ are singleton sets. We say that $X$ is 0-dimensional if $\operatorname{clop}(X)$ forms a basis for $X$. When $X$ is a compact Hausdorff space, the conditions of being totally disconnected and being 0-dimensional are equivalent. A compact Hausdorff space satisfying these conditions is called a Stone space.
We note that Gelfand duality implies that, for compact Hausdorff spaces $X$ and $Y$, $X$ and $Y$ are homeomorphic if and only if $C(X)\cong C(Y)$ as $\mathrm{C}^{\ast}$}-algebras. Similarly, Stone duality (see [Farah:2019, §1.3.1]) implies that, for Stone spaces $X$ and $Y$, $X$ and $Y$ are homeomorphic if and only if $\operatorname{clop}(X)\cong\operatorname{clop}(Y)$ as Boolean algebras. Moreover, a compact Hausdorff space $X$ is a Stone space if and only if $C(X)$ is an abelian $\mathrm{C}^{\ast}$}-algebra of real rank zero.
Lemma 6.4 ([EAGLE:2015, Lemma 5.7]).
Let $X$ be a compact Hausdorff space, and $\mathcal{U}$ be an ultrafilter. Then $C(X)^{\mathcal{U}}\cong C(\sum_{\mathcal{U}}X)$, and $\operatorname{clop}(X)^{\mathcal{U}}\cong\operatorname{clop}(\sum_{\mathcal{U}}X)$ where $\sum_{\mathcal{U}}X$ is the ultracopower of $X$ and $\operatorname{clop}(X)^{\mathcal{U}}$ is the classical model-theoretic ultrapower of Boolean algebra $\operatorname{clop}(X)$.
For a topological space $X$, the weight of $X$, denoted by $w(X)$, is the smallest cardinality of a basis for $X$.
Lemma 6.5 ([FM:2021, Lemma 1.2]).
Suppose that $X$ is an infinite Hausdorff space. Then, the density character of $C(X)$ is equal to $w(X)$.
Therefore, by applying section 5 to $M=C(2^{\omega})$, we obtain the following result:
Corollary 6.6.
[…]
  1. (1)
    $w(X),w(Y)\leq{\mathfrak{c}}$.
  2. (2)
    There exist a set $I$ and non-principal ultrafilters $\mathcal{U},\mathcal{V}\in\beta I\setminus I$ such that $\sum_{\mathcal{U}}X$ and $\sum_{\mathcal{V}}Y$ are homeomorphic.
  3. (3)
    For any $\mathcal{U},\mathcal{V}\in\beta\omega\setminus\omega$, $\sum_{\mathcal{U}}X$ and $\sum_{\mathcal{V}}Y$ are not homeomorphic.
Proof.
Suppose that CH fails. By section 5 with $M=C(2^{\omega})$, there exist $A,B\equiv M$ such that $\chi(A),\chi(B)\leq{\mathfrak{c}}$ and $A^{\mathcal{U}}\not\cong B^{\mathcal{V}}$ for all $\mathcal{U},\mathcal{V}\in\beta\omega\setminus\omega$. By section 3.1, there exist a set $I$ and non-principal ultrafilters $\mathcal{U},\mathcal{V}\in\beta I\setminus I$ such that $A^{\mathcal{U}}\cong B^{\mathcal{V}}$.
Note that $M$ is a unital abelian $\mathrm{C}^{\ast}$}-algebra of real rank zero. Since the class of real rank zero $\mathrm{C}^{\ast}$}-algebras is elementary (see [MToCA, Example 2.4.2]), $A$ and $B$ are also unital abelian $\mathrm{C}^{\ast}$}-algebras of real rank zero. Thus, there exist Stone spaces $X$ and $Y$ such that $A\cong C(X)$ and $B\cong C(Y)$.
By section 6, we have $w(X),w(Y)\leq{\mathfrak{c}}$, that is, condition (1) holds. Moreover, conditions (2) and (3) also hold for $X$ and $Y$, by the properties of $A$ and $B$ and by section 6. ∎
Thus, section 6 for abelian $\mathrm{C}^{\ast}$}-algebras can be reformulated as the following question about topological spaces:
Question 6.7.
[…]
  1. (1)
    $w(X),w(Y)\leq\aleph_{0}$, that is, $X$ and $Y$ are second countable.
  2. (2)
    There exist a set $I$ and non-principal ultrafilters $\mathcal{U},\mathcal{V}\in\beta I\setminus I$ such that $\sum_{\mathcal{U}}X$ and $\sum_{\mathcal{V}}Y$ are homeomorphic.
  3. (3)
    For any $\mathcal{U},\mathcal{V}\in\beta\omega\setminus\omega$, $\sum_{\mathcal{U}}X$ and $\sum_{\mathcal{V}}Y$ are not homeomorphic.
Remark 6.8.
By the Riesz-Markov-Kakutani theorem, for a compact Hausdorff space $X$, $C(X)$ is separable if and only if $X$ is metrizable. Hence, in section 6 the requirement that $X$ and $Y$ are second countable can be replaced with the requirement that $X$ and $Y$ are metrizable.